### v1

# WaterSeeker: Efficient Detection of Watermarked Segments in Large Documents

Abstract
Watermarking algorithms for large language models (LLMs) have attained high accuracy in detecting LLM-generated text. However, existing methods primarily focus on distinguishing fully watermarked text from non-watermarked text, overlooking real-world scenarios where LLMs generate only small sections within large documents. In this scenario, balancing time complexity and detection performance poses significant challenges. This paper presents WaterSeeker, a novel approach to efficiently detect and locate watermarked segments amid extensive natural text. It first applies an efficient anomaly extraction method to preliminarily locate suspicious watermarked regions. Following this, it conducts a local traversal and performs full-text detection for more precise verification. Theoretical analysis and experimental results demonstrate that WaterSeeker achieves a superior balance between detection accuracy and computational efficiency. Moreover, WaterSeeker’s localization ability supports the development of interpretable AI detection systems. This work pioneers a new direction in watermarked segment detection, facilitating more reliable AI-generated content identification.
1 Introduction
As the quality of text generated by large language models (LLMs) improves, these models address various practical challenges while also raising concerns such as misinformation (Liu et al., 2023c; Chen & Shu, 2024) and copyright infringement (Rillig et al., 2023). To tackle these issues, LLM watermarking technology has emerged, embedding specific information (watermarks) during text generation for accurate detection through specialized algorithms. Existing watermarking methods mainly focus on distinguishing fully watermarked text from non-watermarked content (Kirchenbauer et al., 2023; Zhao et al., 2024; Liu et al., 2024a; Lu et al., 2024; Aaronson & Kirchner, 2022; Kuditipudi et al., 2024). However, in real-world scenarios, LLMs often generate only brief segments within longer documents, making it essential to detect these watermarked sections among large amounts of non-watermarked text.
Most previous algorithms relied on computing statistics across the entire document for detection, which we refer to as full-text detection methods. For instance, some approaches categorize vocabulary into green and red lists, applying a bias towards the selection of green tokens during generation and using metrics like the z-score to measure their proportion throughout the entire text (Kirchenbauer et al., 2023; Zhao et al., 2024; Liu et al., 2024a). Other methods utilize pseudo-random numbers to guide token sampling, calculating correlations across the entire document during detection (Aaronson & Kirchner, 2022; Christ et al., 2024; Kuditipudi et al., 2024). Ultimately, the presence of a watermark is determined by whether these statistics exceed a specified threshold.
However, these full-text detection methods struggle when small amounts of watermarked text are mixed with large volumes of natural text due to dilution effects, as illustrated in Figure 1(a). To the best of our knowledge, the only existing work addressing this scenario is the WinMax (Kirchenbauer et al., 2024) algorithm, which examines all possible window sizes and traverses the full text for each size. However, this method suffers from high time complexity. Additionally, our self-constructed baseline, the Fixed-Length Sliding Window (FLSW) method, faces the issue of watermark strength dilution due to inflexible window sizes, which negatively impacts detection performance.
(a) Full-text detection methods often struggle in watermarked segment detection scenarios due to the dilution effect.
(b) An illustration of the WaterSeeker algorithm, using “first locate, then detect” strategy to balance time complexity and detection performance.
Figure 1: An overview of watermark detection scenarios and the proposed WaterSeeker algorithm.
To address these issues, we propose a novel watermark detection method named WaterSeeker, which is adaptable to various watermarking techniques. This method achieves a superior balance between time complexity and detection performance when identifying watermarked segments within large documents. The core concept of WaterSeeker is “first locate, then detect”. As illustrated in Figure 1(b), it initially employs a low-complexity anomaly points extraction algorithm to identify suspected watermark regions. This process narrows the detection target from a long text down to a small segment that encompasses the ground truth segment. Next, a local traversal is performed on the localization result. Within each window, full-text watermark detection is conducted, and the highest confidence result is compared to a threshold for the final determination. Theoretical analysis indicates that this coarse-to-fine process has the potential to achieve optimal detection performance while maintaining the lowest possible complexity for solving this problem.
In the experiment, we compared the effectiveness of WaterSeeker and the baselines in detecting watermarked segments of random lengths inserted into large documents, as well as their time complexity. The results show that WaterSeeker significantly surpasses the baseline methods in effectively balancing time complexity and detection performance. Additionally, more detailed analysis includes an adaptability assessment of WaterSeeker under varying watermark strengths and segment lengths, as well as an evaluation of its two-stage utility via ablation study. Furthermore, we discuss WaterSeeker’s contribution to building more interpretable AI detection systems.
In summary, the contributions of this work are as follows:
  • We comprehensively define the new scenario of detecting watermarked segments in large documents, specifying the inputs and outputs of the algorithms designed to address this problem, as well as the evaluation metrics and the construction process for the test dataset.
  • We propose WaterSeeker, a novel and general watermark detection method that effectively identifies watermarked segments in large documents, addressing the challenges posed by dilution effects in non-watermarked text.
  • WaterSeeker achieves a superior balance between time complexity and detection effectiveness, with strong adaptability to varying watermark strengths and lengths, significantly outperforming existing baselines.
  • WaterSeeker not only provides accurate detection results but also offers precise localization outcomes, contributing to the development of more interpretable AI detection systems.
2 Related Work
Watermarking for large language models (LLMs) embeds subtle, algorithmically detectable signals in outputs to identify LLM-generated text, making it crucial for mitigating potential misuse of LLMs. Currently, mainstream watermarking methods involve intervening in the inference phase of LLMs by modifying the logits output or influencing the token sampling process to produce watermarked text (Liu et al., 2023b; Pan et al., 2024; Liu et al., 2024b).
The KGW method (Kirchenbauer et al., 2023) modifies logits by categorizing the vocabulary into green and red lists, applying a bias towards green tokens during text generation. The bias value is typically determined by the parameter $\delta$, which reflects the watermark strength. For detection, it first labels each token as red or green and calculates the z-score of the entire text to reflect the proportion of green tokens; if this proportion exceeds a specified threshold, the text is classified as watermarked. In subsequent work, numerous enhancements have been suggested to improve text quality (Hu et al., 2024; Wu et al., 2023; Takezawa et al., 2023), increase information capacity (Wang et al., 2024; Yoo et al., 2023; Fernandez et al., 2023), boost robustness against watermark removal attacks (Zhao et al., 2024; Liu et al., 2024a; Ren et al., 2024; He et al., 2024), adapt to low-entropy conditions (Lee et al., 2023; Lu et al., 2024), and facilitate public detection (Liu et al., 2023a; Fairoze et al., 2023).
On the other hand, the Aar method (Aaronson & Kirchner, 2022) proposes utilizing pseudo-random sequences to guide the token sampling process. During each token sampling step, a pseudo-random vector $u\sim\text{Uniform}([0,1])^{|V|}$ is generated based on the previously sampled tokens. Assuming that the probability vector provided by the LLM is $p$, the method performs exponential sampling to select the token $i$ that maximizes $u_{i}^{1/p_{i}}$. Here, the watermark strength is reflected by a sampling temperature, which affects the strictness of the exponential sampling. For detection, it also employs global statistics: it sums the correlation values of each token with the pseudo-random vector and then performs gamma transformation to derive the detection confidence. Building on this, Kuditipudi et al. (2024) introduced the use of edit distance to calculate correlation, while Hou et al. (2024a) and Hou et al. (2024b) incorporated contextual semantic information to guide the sampling process, thereby enhancing detection robustness.
Despite the high accuracy of watermarking algorithms for detecting LLM-generated text, previous studies have primarily focused on distinguishing between fully watermarked and non-watermarked text, overlooking the possibility that LLMs may only generate small segments within large documents. In such cases, watermark detection algorithms based on full-text statistics can fail due to the dilution effect. A few studies have mentioned copy-paste attack (Kirchenbauer et al., 2024; Yoo et al., 2023; Wang et al., 2024), where a certain proportion of watermarked text is mixed into non-watermarked content, resembling our scenario. Yoo et al. (2023) and Wang et al. (2024) tested the robustness of their methods against copy-paste attacks by mixing 10% to 50% watermarked text into non-watermarked text. However, since they did not specifically design detection mechanisms for this scenario, their results indicated a lack of robustness against such attack.
In existing studies, WinMax (Kirchenbauer et al., 2024) is the only watermark detection algorithm specifically designed for this scenario. It examines all possible window sizes and traverses the entire text for each size to identify the maximum local score for threshold comparison. While this method shows some effectiveness, its high time complexity renders it impractical for real-world applications. Therefore, to achieve effective detection of watermarked segments inserted into large documents while maintaining controllable time complexity, we propose WaterSeeker, a novel and general method that uses “first locate then detect” strategy. We validate its superiority through comparisons with full-text detection, WinMax, and a self-constructed fixed-length sliding window detection method.
3 Problem Formulation
3.1 Definition
The problem of detecting watermarked segments in a long document is defined as follows: Let $N$ be a long text with a randomly inserted watermarked segment of length $L$ (where $L$ is a random variable). We denote the starting index of the watermarked segment in $N$ as $s$ and the ending index as $e$, such that: $L=e-s+1$. The objective is to determine the presence and location of a watermarked segment in $N$. This can be framed as a binary classification problem with additional localization, where the detection algorithm outputs:
[…]
Here, ‘has_watermark’ is a boolean value indicating the presence of a watermark, and ‘indices’ is a list of pairs $(s^{\prime},e^{\prime})$ representing the start and end indices of detected watermarked segments.
3.2 Evaluation
A watermark is considered successfully detected if:
[…]
  • There exists at least one detected segment $(s^{\prime},e^{\prime})$ such that the Intersection over Union (IoU) with the ground truth segment $(s,e)$ exceeds a specified threshold $\theta$:
    $\text{IoU}=\cfrac{L_{\text{intersection}}}{L_{\text{union}}}>\theta.$ (1)
Based on this evaluation, we will report the following metrics of the binary classification result: False Positive Rate (FPR), False Negative Rate (FNR), and F1 Score. Additionally, we will include the average IoU between the detected segments and ground truth segments as a supplementary metric, further demonstrating the accuracy of the localization.
4 Baseline Methods
4.1 Full-text Detection
[…]
4.2 WinMax
WinMax (Kirchenbauer et al., 2024) is a watermark detection algorithm proposed on the basis of KGW to address copy-paste attack. It involves iterating through all possible window sizes, and for each window size, the entire text is traversed to calculate the z-score for each local window, taking the maximum z-score and comparing it against a specified threshold. In practical applications, a minimum window size $w_{min}$ and a maximum window size $w_{max}$ are typically set to control the number of iterations during the traversal. The detection process can be described by the following formula:
$z_{\text{win-max}}=\max_{w\in[w_{min},w_{max}]}\left(\max_{i}\left(\frac{|s|_{G,i}-\gamma w}{\sqrt{\gamma(1-\gamma)w}}\right)\right),$ (2)
where $w$ is the length of the local window, and $|s|_{G,i}$ represents the count of green tokens within the $i$-th local window of length $w$. For other watermarking algorithms, such as Aar, a similar approach can also be implemented. The time complexity of the WinMax detection method is relatively high, costing $O((w_{max}-w_{min})\times N)$, where $N$ is the total length of the text. Due to the uncertain length of the watermarked segment inserted into the document, if there is a large difference between $w_{max}$ and $w_{min}$, the worst-case complexity can reach $O(N^{2})$.
4.3 Fix-Length Sliding Window
Fix-Length Sliding Window (FLSW) detection method is a self-constructed, straightforward approach to address the watermark detection problem in this scenario. It involves using a fixed-length window to traverse the entire text, calculating statistics within the local window. If the statistics within a local window exceed a specified threshold, the variable ‘has_watermarked’ is set to true. For localization, during the traversal, each pair of indices $(s,e)$ that meets the condition is recorded. Finally, a contiguous method with a certain tolerance is employed to provide the localization results. The pseudocode for the aforementioned baselines can be found in Appendix A.
Although the concept of FLSW is relatively simple and straightforward, the fixed-length nature makes it unable to adapt to varying watermark lengths. This limitation can result in dilution effects when the window size is either too large or too small. In Section 5.1, we will analyze this dilution effect in detail using KGW (Kirchenbauer et al., 2023) and Aar (Aaronson & Kirchner, 2022) as examples, providing theoretical support for WaterSeeker.
5 Proposed Method: WaterSeeker
5.1 Theoretical Foundation: Gold Index is the Best
In this section, we will conduct a theoretical analysis to demonstrate that using the actual start and end indices (the gold index) for watermark detection yields the highest expected detection rate, using KGW and Aar as examples. As mentioned in Section 2, most watermarking algorithms are slight modifications of these two methods, maintaining the fundamental principles of statistical calculation. Therefore, analyzing these two approaches provides a broad applicability.
For the KGW method, assuming $\gamma_{1}>\gamma$ is the proportion of green tokens in the watermarked part. Let’s analyze the effect of window size $W$ on this statistic:
  • When $W<L$ (window size smaller than watermark length), we have $E[|s|_{G}]=W\gamma_{1}$. Therefore,
    $E[z_{W}]=\frac{W\gamma_{1}-\gamma W}{\sqrt{\gamma(1-\gamma)W}}=\sqrt{W}\cdot\frac{\gamma_{1}-\gamma}{\sqrt{\gamma(1-\gamma)}},$ (3)
    where the expectation of z-score is proportional to $\sqrt{W}$.
  • When $W>L$ (window size larger than watermark length), we have $E[|s|_{G}]=L\gamma_{1}+(W-L)\gamma$. Therefore
    $E[z_{W}]=\frac{L\gamma_{1}+(W-L)\gamma-\gamma W}{\sqrt{\gamma(1-\gamma)W}}=\frac{L(\gamma_{1}-\gamma)}{\sqrt{\gamma(1-\gamma)W}},$ (4)
    where the expectation of z-score is proportional to $1/\sqrt{W}$, decreasing as $W$ increases.
From this, we can conclude that when $W=L$, the z-score reaches its maximum.
During detection, we aim for a higher z-score for positive cases while setting an appropriate threshold to balance the false positive rate. Next, we will analyze the constraints on the z-threshold $z^{*}$ when the false positive rate within the specified window is set to be lower than a target value $\alpha$.
We start with the binomial distribution $B(W,\gamma)$ that describes the number of green tokens in a window of size $W$. For large sample sizes (typically $W>50$), we can apply the Central Limit Theorem to approximate this binomial distribution with a normal distribution $N(W\gamma,W\gamma(1-\gamma))$. We then standardize this to a z-score, which follows the standard normal distribution $N(0,1)$. The false positive rate alpha represents the area in the right tail of this distribution beyond $z^{*}$. Mathematically, this is expressed as $\alpha=1-\Phi(z^{*})$, where $\Phi$ is the cumulative distribution function (CDF) of the standard normal distribution. Solving for $z^{*}$, we get:
$z^{*}=\Phi^{(-1)}(1-\alpha),$ (5)
which is a constant value for different $W$.
For the Aar method, the p-value is calculated by applying a gamma transformation to the sum of correlation values of all the tokens:
[…]
Due to the complexity of the Gamma-Transformation, the detailed proof can be found in Appendix B. Based on the above proof, we simulated the expected score of the statistic and the corresponding threshold as they vary with $W$ for $L=200$ using real data, as shown in Figure 2. It demonstrates that when the window size corresponds to the gold index, the highest detection rate can be achieved while maintaining an acceptable FPR. Therefore, the watermark dilution effect refers to the situation where the window used for detection is not optimal, resulting in the computed statistics representing watermark strength not reaching their peak or minimum, which in turn affects detection performance.
(a) KGW, $\alpha=10^{-6},\gamma=0.5,\gamma_{1}=0.75$
(b) Aar, $\alpha=10^{-6},\mu=1.0,\mu_{1}=1.6$
Figure 2: Expected score of the statistics and the corresponding threshold across varying $W$.
5.2 Suspicious Region Localization
Based on the theoretical analysis presented in Section 5.1, the design concept of WaterSeeker employs a coarse-to-fine process to gradually approximate the gold index, thereby achieving high detection performance. In the first step (coarse step), WaterSeeker uses a localization algorithm to identify suspicious watermarked regions. This process narrows the detection target to a small segment that encompasses the ground truth segment, while ensuring relatively small deviation from the gold index. Below is the detailed process of the localization algorithm:
  1. 1.
    Score List Computation: We utilize a small sliding window (i.e., $W=50$) to efficiently traverse the entire text. For each position of the window, we compute a statistical measure that reflects the average watermark intensity. For example, in the KGW method, this statistic is represented as $|s|_{G}/W$, indicating the density of green tokens inside the window. Meanwhile, in the Aar method, it is calculated as $\cfrac{\sum\log(1/(1-u))}{W}$, representing the average correlation value between tokens and the pseudo-random sequence within the window. This results in a score list $s$ of length $N-W+1$, where $s_{i}$ represents the average watermark intensity of $n_{i}$ to $n_{i}+W$.
  2. 2.
    Anomaly Extraction: In this step, we design an anomaly extraction algorithm to identify outliers from the score list. We calculate the mean of the score list, the top-k mean, and the standard deviation, denoted as $s_{\text{mean}}$, $s_{\text{top-k-mean}}$, and $s_{\text{std}}$, respectively. The extracted outliers are the points that satisfy the following condition, where $\theta_{1}$ and $\theta_{2}$ are hyperparameters:
    $\text{score}>s_{\text{mean}}+\max((s_{\text{top-k-mean}}-s_{\text{mean}})\cdot\theta_{1},s_{\text{std}}\cdot\theta_{2}).$ (7)
    In samples containing a watermarked segment, the first term in the max expression becomes significant. Introducing the top-k mean allows for adaptive adjustment of the deviation from the mean based on different watermark strengths. Conversely, in non-watermarked samples, the difference between the top-k mean and the mean is relatively small. In this case, to reduce false positives, a point must deviate from the mean by at least $s_{\text{std}}\cdot\theta_{2}$ to be considered an anomaly.
  3. 3.
    Fragment Connection: In this step, we use a connecting method with a certain tolerance to link nearby outliers, then filter out segments that are too short, returning a list of indices.
The scores within watermarked segments tend to be relatively stable and close to the top-k mean. Concurrently, the scores corresponding to segments outside the watermark also remain stable and approximately equal to the overall mean (because a large volume of natural text dilutes the overall watermark strength). As a result of these characteristics, when we use $(s_{\text{top-k-mean}}-s_{\text{mean}})\cdot\theta_{1}\quad(\theta_{1}<1)$ as a criterion for distinction, the extracted abnormal segment’s start and end points ($s^{\prime}$ and $e^{\prime}$) generally satisfy the following conditions: $s^{\prime}\in(s-W,s)$ and $e^{\prime}\in(e,e+W)$. This ensures that the extracted suspicious watermarked regions are likely to cover the actual segments while keeping the starting and ending deviations within a window size.
Algorithm 1 WaterSeeker Algorithm
1:procedure SuspiciousRegionLocalization(tokens, $W$, $k$, $\theta_{1}$, $\theta_{2}$)
2:     Compute score list using sliding window
3:     Calculate $s_{\text{mean}}$, $s_{\text{top-k-mean}}$, and $s_{\text{std}}$
4:     Detect anomalies using threshold:
5:     $s_{\text{mean}}+\max((s_{\text{top-k-mean}}-s_{\text{mean}})\cdot\theta_{1},s_{\text{std}}\cdot\theta_{2})$
6:     Connect nearby outliers and filter short segments
7:     return filteredSegments
8:end procedure
9:procedure LocalTraverseDetection(tokens, suspiciousRegions, $W$, threshold)
10:     hasWatermark $\leftarrow$ False
11:     indices $\leftarrow$ [ ]
12:     for $(s^{\prime},e^{\prime})$ in suspiciousRegions do
13:         maxScore = $\infty$- ∞, maxIndice = ()
14:         for $s\in[s^{\prime},s^{\prime}+W)$ and $e\in(e^{\prime}-W,e^{\prime}]$ do
15:              if DetectWatermark(tokens[$s:e$])$>$ maxScore then
16:                  maxScore = DetectWatermark(tokens[$s:e$])
17:                  maxIndice = $(s,e)$
18:              end if
19:         end for
20:         if maxScore $>$ threshold then
21:              hasWatermark $\leftarrow$ True
22:              indices.append(maxIndice)
23:         end if
24:     end for
25:     return hasWatermark, indices
26:end procedure
27:procedure WaterSeeker(tokens, $W$, $k$, $\theta_{1}$, $\theta_{2}$, threshold)
28:     suspiciousRegions $\leftarrow$ SuspiciousRegionLocalization(tokens, $W$, $k$, $\theta_{1}$, $\theta_{2}$)
29:     hasWatermark, indices $\leftarrow$ LocalTraverseDetection(tokens, suspiciousRegions, W, threshold)
30:     return hasWatermark, indices
31:end procedure
5.3 Local Traverse Detection
After obtaining the localization results from the first step, the second phase (fine step), local traverse detection, conducts a more detailed verification. For each $(s^{\prime},e^{\prime})$ pair in the localization results, traverse inwards through the segments where the start point falls within $[s^{\prime},s^{\prime}+W)$ and the end point within $(e^{\prime}-W,e^{\prime}]$. Based on the previous analysis of the ranges of $s^{\prime}$ and $e^{\prime}$, this traversal has a high probability of reaching the gold index. Perform full-text detection sequentially on these segments, and select the statistic with the most significance (e.g., for KGW, choose the segment with the largest z-score, and for Aar, the segment with the smallest p-value) to compare with the threshold. The pseudocode for the entire WaterSeeker algorithm can be found in Algorithm 1.
5.4 Time Complexity Analysis
Time Complexity Analysis of WaterSeeker:
  1. 1.
    Suspicious Region Localization: $O(N)$. This step involves three sub-steps: score list computation, anomaly detection, and fragment connection. Each of these sub-steps has a time complexity of $O(N)$, where $N$ is the length of the text.
  2. 2.
    Local Traverse Detection: $O(W^{2})$. For each suspicious region identified, this step examines $W^{2}$ windows, where $W$ is the window size.
The total time complexity of WaterSeeker is $O(N+W^{2})$. In practice, $W^{2}$ is typically maintained at a value lower than $T$, as even a slightly larger window (i.e., $W=50$) is sufficient to provide a relatively smooth and low-noise representation of the surrounding watermark intensity. Therefore, the dominant term in most cases is $O(N)$, making the overall time complexity of WaterSeeker $O(N)$.
Lower Bound Complexity Analysis for the Problem: To detect watermarked segments in a long text, any algorithm must examine each token in the text at least once. This necessitates at least one full pass through the text, establishing a lower bound of $\Omega(N)$ for the time complexity of this problem.
In conclusion, the WaterSeeker algorithm, with its $O(N)$ complexity in typical cases, achieves a time complexity that matches the theoretical lower bound of the problem, indicating that it is asymptotically optimal for the task of detecting watermarked segments in long texts.
6 Experiment
6.1 Experiment Settings
Watermarking Methods and Language Models: We selected two representative watermarking algorithms, KGW (Kirchenbauer et al., 2023) and Aar (Aaronson & Kirchner, 2022), each with three strength levels. For KGW, the strength is determined by parameter $\delta$, with values 2.0, 1.5, and 1.0 (from strong to weak). For Aar, the strength is controlled by parameter temperature, with values 0.3, 0.2, and 0.1 (from strong to weak). We used Llama-2-7b (Touvron et al., 2023) and Mistral-7b (Jiang et al., 2024) as generation models.
Dataset Construction: Following previous studies, the C4 dataset (Raffel et al., 2020) was used for prompts, with the first 30 tokens of each entry serving as input. Watermarked segments of random length (100 to 400 tokens) were generated using randomly selected watermark strengths. For positive examples, one such segment was randomly inserted into each 10,000-token Wikipedia passage (Foundation, ). Negative examples consist of unmodified 10,000-token Wikipedia corpus. Based on this procedure, four datasets were created, each containing 300 positive and 300 negative examples: KGW-llama, KGW-mistral, Aar-llama, and Aar-mistral. These datasets employ different combinations of watermarking methods and LLMs. The specific distributions of watermark strengths and lengths are detailed in Appendix D.
It is important to note that watermarked segments with higher strength and longer length are inherently easier to detect and demonstrate better robustness against the dilution effect. Therefore, to showcase the superiority and adaptability of WaterSeeker, each dataset incorporates watermarked segments of varying strengths and lengths.
Baselines: As introduced in Section 4, Full-text Detection and WinMax (Kirchenbauer et al., 2024) are chosen, as well as Fix-Length Sliding Window method with $W$ of 100, 200, 300 and 400.
Hyper-parameters: The parameters related to WaterSeeker are as follows: $W=50,k=20$, with a tolerance for fragment connection set to 100. The parameter $\theta_{1}$ is set to 0.5, and $\theta_{2}$ is set to 1.5 for both algorithms. For full-text detection in baselines and during local traversal in WaterSeeker, the specific threshold selection is detailed in Appendix C. Notably, careful threshold selection is crucial for maintaining an acceptable false positive rate, as traversing long texts is prone to accumulating false positives. Evaluation-related parameters: $\theta$ mentioned in Section 3.2 is set to 0.5.
6.2 Detection Performance Analysis
Table 1: We compared the detection performance of WaterSeeker with other methods, including Plain Detect, WinMax (Kirchenbauer et al., 2024), and FLSW. The reported metrics include False Positive Rate (FPR), False Negative Rate (FNR), F1 score, and the average Intersection over Union (IoU) between detected segments and ground truth segments in positive samples.
Model Method KGW Aar
FPR FNR F1 IoU FPR FNR F1 IoU
Llama-2-7b Full-text Detection 0.017 1.000 0.000 0.001 0.017 1.000 0.000 0.001
WinMax 0.033 0.257 0.837 0.671 0.010 0.303 0.816 0.629
FLSW-100 0.017 0.480 0.677 0.434 0.000 0.473 0.690 0.430
FLSW-200 0.010 0.413 0.734 0.452 0.003 0.543 0.626 0.371
FLSW-300 0.010 0.563 0.604 0.359 0.003 0.787 0.351 0.294
FLSW-400 0.007 0.807 0.322 0.272 0.007 0.893 0.192 0.222
WaterSeeker(Ours) 0.030 0.270 0.830 0.653 0.007 0.330 0.799 0.602
Mistral-7b Full-text Detection 0.030 1.000 0.000 0.001 0.017 1.000 0.000 0.002
WinMax 0.060 0.147 0.892 0.745 0.043 0.453 0.688 0.488
FLSW-100 0.010 0.397 0.748 0.509 0.013 0.577 0.589 0.332
FLSW-200 0.013 0.343 0.786 0.510 0.017 0.657 0.505 0.295
FLSW-300 0.023 0.573 0.589 0.389 0.017 0.807 0.320 0.233
FLSW-400 0.027 0.840 0.271 0.289 0.017 0.933 0.123 0.171
WaterSeeker(Ours) 0.047 0.173 0.883 0.709 0.037 0.483 0.665 0.457
In Table 1, we present the detection performance of WaterSeeker alongside various baseline algorithms, evaluated on four datasets. The results demonstrate that WaterSeeker achieves lower FNR and higher F1 score, significantly surpassing other watermark detection algorithms, including Full-text Detection, FLSW-100, FLSW-200, FLSW-300, and FLSW-400, while being comparable to WinMax. Based on the principle that WinMax is guaranteed to reach the gold index by evaluating all possible windows, it represents the upper bound of detection performance. However, subsequent experiments show that WinMax has extremely high time complexity (Section 6.4). Furthermore, the IoU results in the table reveal that WaterSeeker possesses a significantly higher localization capability compared to the baselines, again comparable to WinMax.
6.3 Adaptability Analysis
While the four datasets used in the main experiment incorporated a mix of watermarked segments with varying strengths and lengths, this section presents a more granular comparison of the detection capabilities of WaterSeeker and the FLSW method (FLSW-100, FLSW-200, FLSW-300, and FLSW-400) across specific watermark strengths and lengths. As shown in Figure 3, the performance of different detection algorithms exhibits a consistent trend: watermarked segments with higher strengths and longer lengths are more easily detected. Moreover, it is evident that FLSW performs well within a length range close to its window size, but shows poor adaptability in other ranges. This limitation arises from its fixed-length nature, which leads to dilution effects when detecting longer or shorter watermarked segments, thereby reducing detection performance. In contrast, WaterSeeker exhibits good adaptability across different length ranges by using anomaly extraction techniques. Additionally, it incorporates a top-k score mechanism to adapt to varying watermark strengths.
(a) KGW, $\delta=2.0$
(b) KGW, $\delta=1.5$
(c) KGW, $\delta=1.0$
(d) Aar, temp=0.3
(e) Aar, temp=0.2
(f) Aar, temp=0.1
Figure 3: This figure compares the detection performance of WaterSeeker and FLSW under varying watermark lengths and strengths, using Llama-2-7b as the generation model.
6.4 Time Complexity Analysis
Table 2: Average detection time per sample for various detection methods ($T\approx 10,000$, Unit: s).
Full-text Detection WinMax FLSW-100 FLSW-200 FLSW-300 FLSW-400 WaterSeeker
KGW 1.70 14.16 1.76 1.76 1.79 1.80 1.75
Aar 0.54 2733.78 1.62 1.65 1.66 1.64 1.68
Table 2 lists the average time spent per sample for WaterSeeker and other baseline algorithms during detection. The results show that the time taken by WaterSeeker is comparable to that of full-text detection and FLSW, and significantly lower than the time required by WinMax. Sections 4.2 and 5.4 analyze the theoretical time complexity of WinMax and WaterSeeker, respectively. This section will further delve into a more in-depth analysis based on experimental results.
In this experimental setting, let the time required to compute a score in the score list be $t_{1}$, and the time required to compute the watermark detection statistic within a local window be $t_{2}$. The time expenditure for the WinMax [^1] The WinMax algorithm, in practical implementation, also requires computing the score list for the entire text first, then traversing the windows to calculate the detection statistic. algorithm to detect a sample can be expressed as $N\cdot t_{1}+N\cdot(W_{max}-W_{min})\cdot t_{2}$. For WaterSeeker, the first step of localization consists of three sub-steps: score list computation, anomaly detection, and fragment connection. Each of these sub-steps has a complexity of $O(N)$. If we denote the time required for the latter two sub-steps as $t_{3}\cdot N$, then the total time for the first step is $N\cdot t_{1}+N\cdot t_{3}$. The second step, local traverse detection, requires $W^{2}\cdot t_{2}$ time. Consequently, the total time expenditure for the entire process can be expressed as $N\cdot t_{1}+N\cdot t_{3}+W^{2}\cdot t_{2}$.
For the Aar (Aaronson & Kirchner, 2022) watermarking method, calculating the watermark detection statistic involves computing a GammaTransform function, which is a relatively complex computation, resulting in a larger value for $t_{2}$. In this case, the time difference between WinMax and WaterSeeker will become significant due to the large disparity in the coefficients preceding $t_{2}$. In this experiment, with $T=10,000,W_{max}=400,W_{min}=100,W^{2}=2,500$, the coefficient difference amounts to a factor of 1,200. For the KGW (Kirchenbauer et al., 2023) watermarking method, calculating the z-score within a local window is a relatively straightforward computation. In this case, the time difference between WinMax and WaterSeeker will be reduced.
Through the above analysis, it is clear that WaterSeeker is a general watermark detection method with consistently low time expenditure across various watermarking algorithms. In contrast, WinMax exhibits extremely high time costs in situations where calculating the watermark detection statistic is complex, making it challenging for practical use.
6.5 Ablation Study
Table 3: This table presents the contributions of the first step of WaterSeeker: Suspicious Segment Localization. It lists the average coverage of localization results relative to the ground truth segments under various watermark algorithms and watermark strengths, as well as the average offsets of the detected start and end indices.
Metrics KGW Aar
$\delta=2.0$ $\delta=1.5$ $\delta=1.0$ temp=0.3 temp=0.2 temp=0.1
Average Coverage 0.992 0.984 0.953 0.996 0.991 0.977
Average Offset (Start) 20.799 30.057 13.992 22.559 30.546 20.624
Average Offset (End) 25.985 25.299 28.030 29.459 33.059 30.304
In this section, we will analyze the effectiveness of the two stages of WaterSeeker through an ablation study. The first stage of WaterSeeker, Suspicious Segment Localization, aims to achieve significant coverage of the ground truth segments while keeping the start and end offsets within a specified window size. This ensures that subsequent local traversals can reach the gold index. Table 3 illustrates the performance of Step 1, showing that the localization results of Suspicious Segment Localization achieve an average coverage level exceeding 0.95 across various watermark algorithms and strengths. Additionally, both the average start offset and average end offset are less than 50, remaining within the designated window size. This indicates that Step 1 fulfills its intended purpose effectively.
Local Traverse Detection performs a localized traversal based on the segments narrowed down in the first step, allowing for more refined verification within the window. Table 4 illustrates the differences in detection outcomes when using Local Traverse compared to directly applying full-text detection with the localization results. It can be found that across different LLMs and watermarking algorithms, Local Traverse consistently enhances the detection performance, making it an indispensable component of WaterSeeker.
Table 4: This table compares the detection performance of using Local Traverse versus not using it across the four datasets.
Model Setting KGW Aar
FPR FNR F1 IoU FPR FNR F1 IoU
Llama-2-7b w. local traverse 0.030 0.270 0.830 0.653 0.007 0.330 0.799 0.602
w/o local traverse 0.010 0.333 0.795 0.565 0.007 0.400 0.747 0.509
Mistral-7b w. local traverse 0.047 0.173 0.883 0.709 0.037 0.483 0.665 0.457
w/o local traverse 0.030 0.267 0.832 0.619 0.027 0.510 0.646 0.400
7 Further Application
Figure 4: An illustration of WaterSeeker applied in AI detection system.
WaterSeeker not only provides accurate detection results, but its localization capabilities also contribute to building a more transparent and interpretable AI detection system. As illustrated in Figure 4, the AI detection system powered by WaterSeeker receives documents for analysis and outputs three components: (1) Detection Result, which indicates whether AI assistance was utilized; (2) Suspicious AI-generated Segments, which highlights segments identified as potentially AI-generated; and (3) AI Ratio, which displays the proportion of content attributed to AI. For example, in the context of academic integrity, highlighting suspicious segments provides actionable evidence for assessments and appeals. Additionally, reporting the AI ratio allows for more flexible establishment of academic misconduct standards.
8 Conclusion
In this work, we defined a new scenario for detecting watermarked segments in large documents and established its evaluation metrics. We identified the limitations of traditional full-text detection methods in this context and proposed a “first locate, then detect” watermark detection algorithm that utilizes a coarse-to-fine strategy. We validated the detection performance and time complexity of our algorithm through a series of analyses and experiments, demonstrating its ability to effectively balance both aspects. For future improvements, we recommend exploring more advanced locating methods based on this concept, which may yield better detection results.
Appendix A Pseudocode of Detection Baselines
Algorithm 2 WinMax Algorithm
1:procedure WinMaxDetection(tokens, $W_{min}$, $W_{max}$, threshold)
2:     hasWatermark $\leftarrow$ False, indices $\leftarrow$ [ ]
3:     maxStat $\leftarrow$ - $\infty$, bestIndex $\leftarrow$ None
4:     for $W$in [$W_{min}$, $W_{max}$do
[…]
6:              stat $\leftarrow$ CalculateStatistics(tokens[$i:i+W$])
7:              if stat $>$ maxStat then
8:                  maxStat $\leftarrow$ stat
9:                  bestIndex $\leftarrow$ $(i,i+W)$
10:              end if
11:         end for
12:     end for
13:     if maxStat $>$ threshold then
14:         hasWatermark $\leftarrow$ True
15:         indices.append(bestIndex)
16:     end if
17:     return hasWatermark, indices
18:end procedure
Algorithm 3 FLSW Algorithm
1:procedure FLSWDetection(tokens, $W$, threshold)
2:     hasWatermark $\leftarrow$ False
3:     indices $\leftarrow$ [ ]
[…]
5:         stat $\leftarrow$ CalculateStatistics(tokens[$i:i+W$])
6:         if stat $>$ threshold then
7:              hasWatermark $\leftarrow$ True
8:              indices.append(($i$, $i+W$))
9:         end if
10:     end for
11:     indices $\leftarrow$ ConnectFragments(indices)
12:     return hasWatermark, indices
13:end procedure
Appendix B Detailed Proof for Aar
For the Aar method, we aim to analyze how the expected p-value for watermarked text varies with changes in the detection window size $W$.
Recall the p-value calculation formula:
[…]
where $S=\sum_{i=1}^{W}\log(\frac{1}{1-u_{i}})$, and $W$ is the window size.
For watermarked tokens, $E[\log(\frac{1}{1-u_{i}})]=\mu_{1}$ , and for non-watermarked tokens, $E[\log(\frac{1}{1-u_{i}})]=\mu_{0}$, where $\mu_{1}>\mu_{0}$.
The expectation of S for different W values is as follows:
  • When $W\leq L$: $E[S]=W\mu_{1}$
  • When $W>L$: $E[S]=L\mu_{1}+(W-L)\mu_{0}$
The expectation of p-value is:
$E[\text{p-value}]=E[\text{GammaTransform}(S,W,\text{loc}=0,\text{scale}=1)]$ (9)
Since GammaTransform is non-linear, we cannot directly substitute $E[S]$. However, we can use Jensen’s inequality to obtain an approximation:
$E[\text{GammaTransform}(S,W,0,1)]\geq\text{GammaTransform}(E[S],W,0,1)$ (10)
Therefore, we can analyze $\text{GammaTransform}(E[S],W,0,1)$ to obtain a lower bound for the expectation of p-value. Define function $f(W)=\text{GammaTransform}(E[S],W,0,1)$:
  • When $W\leq L$: $f(W)=\text{GammaTransform}(W\mu_{1},W,0,1)$
  • When $W>L$: $f(W)=\text{GammaTransform}(L\mu_{1}+(W-L)\mu_{0},W,0,1)$
Analyzing the behavior of $f(W)$:
  • When $W\leq L$, both $E[S]$ and the shape parameter W increase as $W$ increases.
  • When $W$ just exceeds $L$, the growth rate of $E[S]$ suddenly decreases (from $\mu_{1}$ to $\mu_{0}$), while the shape parameter W continues to increase linearly.
Consider the behavior of $f(W)$ near $W=L$:
  • When $W$ increases from $L-\epsilon$ to $L$, $E[S]$ increases by $\epsilon\mu_{1}$, and the shape parameter increases by $\epsilon$.
  • When $W$ increases from $L$ to $L+\epsilon$, $E[S]$ increases by $\epsilon\mu_{0}$, and the shape parameter increases by $\epsilon$.
Since $\mu_{1}>\mu_{0}$, at the point $W=L$, the rate of decrease of $f(W)$ suddenly slows down. This suggests that $f(W)$ is likely to reach its minimum value at $W=L$, or at a point very close to $L$. While this analysis does not strictly prove that the expectation of p-value is minimized exactly at $W=L$, it strongly suggests that the expectation of p-value is likely to reach its minimum value at or very near $W=L$ [^2] For a more rigorous proof, a deeper mathematical analysis of the GammaTransform function or numerical simulations would be necessary.
Similar to the analysis for KGW method, we also need to analyze the constraints on the p-value threshold $p^{*}$ when the false positive rate within the specified window is set to be lower than a target value $\alpha$.
In the Aar method, the test statistic $S$ follows a Gamma distribution with shape parameter $W$ and scale parameter 1. For non-watermarked text, we have: $S\sim\text{Gamma}(W,1)$. The p-value is calculated using the Gamma CDF:
[…]
where GammaCDF is the cumulative distribution function of the Gamma distribution with shape parameter $W$ and scale parameter 1. To achieve a false positive rate of $\alpha$, we need to set a threshold $p^{*}$ such that: $P(\text{p-value}<p^{*})=\alpha$.
Given the definition of p-value, this is equivalent to: $P(1-\text{GammaCDF}(S,W,1)<p^{*})=\alpha$, which can be rewritten as: $P(S>\text{GammaInv}(1-p^{*},W,1))=\alpha$, where GammaInv is the inverse of the Gamma CDF.
Since $S$ follows a $\text{Gamma}(W,1)$ distribution for non-watermarked text, we can express this as:
$1-\text{GammaCDF}(\text{GammaInv}(1-p^{*},W,1),W,1)=\alpha.$ (12)
Solving this equation for $p^{*}$, we get $p^{*}=\alpha$, which is also a constant value for different $W$.
Appendix C Detail of Threshold Selection in Full-text Detection
A key role of threshold selection is to control the false positive rate. In this context, the task involves detecting watermark fragments within long texts, which requires traversing extensive content and can lead to an accumulation of false positives. Therefore, managing the false positive rate within the detection window is crucial in this scenario.
Assuming the false positive rate of event $X$ is $\alpha$, let event $Y$ represent ”event $X$ has at least one false positive in 10,000 independent trials. ” We know that $P(Y)=1-(1-\alpha)^{10000}$. If we want to control $P(Y)$ to be around 0.01, then $\alpha$ should be set to approximately $10^{-6}$. Next, we will explain how to set thresholds for both KGW and Aar to ensure the theoretical false positive rate in window detection is below $10^{-6}$.
For KGW, as analyzed in Section 5.1, when the window size is large, we can approximate using the Central Limit Theorem, resulting in $z^{*}=\Phi^{-1}(1-\alpha)$. When $\alpha=10^{-6}$, this gives $z\approx 4.75$. However, when the window size $W$ is small, the approximation to a normal distribution using the Central Limit Theorem may lead to significant deviations. Therefore, we will use the binomial distribution for precise calculations. $x\sim B(W,\gamma)$ describes the number of green tokens in a window of size $W$ follows a binomial distribution, therefore:
$z=\frac{x-\gamma W}{\sqrt{W\gamma(1-\gamma)}}.$
To find $P(z\geq z^{*})$:
$P(z\geq z^{*})=P\left(\frac{x-\gamma W}{\sqrt{W\gamma(1-\gamma)}}\geq z^{*}\right).$
Expanding this, we have:
$P(z\geq z^{*})=\sum_{k=0}^{W}\binom{W}{k}\gamma^{k}(1-\gamma)^{W-k}\mathbb{I}\left\{\frac{k-\gamma W}{\sqrt{W\gamma(1-\gamma)}}\geq z^{*}\right\}.$
This is the exact expression for $P(z\geq z^{*})$ without any approximations.
We can further simplify:
$P(z\geq z^{*})=\sum_{k=0}^{W}\binom{W}{k}\gamma^{k}(1-\gamma)^{W-k}\mathbb{I}\left\{k\geq\gamma W+z^{*}\sqrt{W(1-\gamma)}\right\}.$
We need to find an appropriate $z^{*}$ such that $P(z\geq z^{*})<\alpha$. This function does not have a direct analytical solution, so we can increment $z^{*}$ in steps of 0.01 until the probability exceeds $\alpha$. The final value of $z^{*}$ is dependent on $W$, and we pre-compute these values during experiments and store them in a dictionary. In experiments, for detected segments with a length of 200 or more, we directly apply the Central Limit Theorem approximation, setting $z=4.75$. For segments shorter than 200, we use the binomial distribution and retrieve the corresponding threshold from the pre-computed dictionary.
For Aar, as analyzed in Appendix B, $p^{*}=\alpha$, therefore set $p^{*}=10^{-6}$ for all lengths.
Appendix D Sample Distribution in the Main Experiment
It can be observed from Figure 5 that all four datasets include samples of watermarked segments with varying intensities and lengths, demonstrating the comprehensiveness and fairness of the dataset construction.
(a) KGW, llama
(b) Aar, llama
(c) KGW, mistral
(d) Aar, mistral
Figure 5: Distribution of sample numbers for the four datasets involved in the main experiment.

### v2

@@
 This work pioneers a new direction in watermarked segment detection, facilitating more reliable AI-generated content identification.
+Our code is available at https://github.com/THU-BPM/WaterSeeker.
 1 Introduction
-As the quality of text generated by large language models (LLMs) improves, these models address various practical challenges while also raising concerns such as misinformation (Liu et al., 2023c; Chen & Shu, 2024) and copyright infringement (Rillig et al., 2023).
+As the quality of text generated by large language models (LLMs) improves, these models address various practical challenges while also raising concerns such as misinformation (Liu et al., 2023b; Chen & Shu, 2024) and copyright infringement (Rillig et al., 2023).
 To tackle these issues, LLM watermarking technology has emerged, embedding specific information (watermarks) during text generation for accurate detection through specialized algorithms.
-Existing watermarking methods mainly focus on distinguishing fully watermarked text from non-watermarked content (Kirchenbauer et al., 2023; Zhao et al., 2024; Liu et al., 2024a; Lu et al., 2024; Aaronson & Kirchner, 2022; Kuditipudi et al., 2024).
+Existing watermarking methods mainly focus on distinguishing fully watermarked text from non-watermarked content (Kirchenbauer et al., 2023; Zhao et al., 2024; Liu et al., 2024b; Lu et al., 2024; Aaronson & Kirchner, 2022; Kuditipudi et al., 2024).
 However, in real-world scenarios, LLMs often generate only brief segments within longer documents, making it essential to detect these watermarked sections among large amounts of non-watermarked text.
 Most previous algorithms relied on computing statistics across the entire document for detection, which we refer to as full-text detection methods.
-For instance, some approaches categorize vocabulary into green and red lists, applying a bias towards the selection of green tokens during generation and using metrics like the z-score to measure their proportion throughout the entire text (Kirchenbauer et al., 2023; Zhao et al., 2024; Liu et al., 2024a).
+For instance, some approaches categorize vocabulary into green and red lists, applying a bias towards the selection of green tokens during generation and using metrics like the z-score to measure their proportion throughout the entire text (Kirchenbauer et al., 2023; Zhao et al., 2024; Liu et al., 2024b).
 Other methods utilize pseudo-random numbers to guide token sampling, calculating correlations across the entire document during detection (Aaronson & Kirchner, 2022; Christ et al., 2024; Kuditipudi et al., 2024).
@@
 Watermarking for large language models (LLMs) embeds subtle, algorithmically detectable signals in outputs to identify LLM-generated text, making it crucial for mitigating potential misuse of LLMs.
-Currently, mainstream watermarking methods involve intervening in the inference phase of LLMs by modifying the logits output or influencing the token sampling process to produce watermarked text (Liu et al., 2023b; Pan et al., 2024; Liu et al., 2024b).
+Currently, mainstream watermarking methods involve intervening in the inference phase of LLMs by modifying the logits output or influencing the token sampling process to produce watermarked text (Liu et al., 2023a; Pan et al., 2024; Liu et al., 2024c).
 The KGW method (Kirchenbauer et al., 2023) modifies logits by categorizing the vocabulary into green and red lists, applying a bias towards green tokens during text generation.
@@
 For detection, it first labels each token as red or green and calculates the z-score of the entire text to reflect the proportion of green tokens; if this proportion exceeds a specified threshold, the text is classified as watermarked.
-In subsequent work, numerous enhancements have been suggested to improve text quality (Hu et al., 2024; Wu et al., 2023; Takezawa et al., 2023), increase information capacity (Wang et al., 2024; Yoo et al., 2023; Fernandez et al., 2023), boost robustness against watermark removal attacks (Zhao et al., 2024; Liu et al., 2024a; Ren et al., 2024; He et al., 2024), adapt to low-entropy conditions (Lee et al., 2023; Lu et al., 2024), and facilitate public detection (Liu et al., 2023a; Fairoze et al., 2023).
+In subsequent work, numerous enhancements have been suggested to improve text quality (Hu et al., 2024; Wu et al., 2023; Takezawa et al., 2023), increase information capacity (Wang et al., 2024; Yoo et al., 2023; Fernandez et al., 2023), boost robustness against watermark removal attacks (Zhao et al., 2024; Liu et al., 2024b; Ren et al., 2024; He et al., 2024), adapt to low-entropy conditions (Lee et al., 2023; Lu et al., 2024), and facilitate public detection (Liu et al., 2024a; Fairoze et al., 2023).
 On the other hand, the Aar method (Aaronson & Kirchner, 2022) proposes utilizing pseudo-random sequences to guide the token sampling process.
@@
 The parameter $\theta_{1}$ is set to 0.5, and $\theta_{2}$ is set to 1.5 for both algorithms.
-For full-text detection in baselines and during local traversal in WaterSeeker, the specific threshold selection is detailed in Appendix C.
+The threshold selection within the specified window is detailed in Appendix C.
 Notably, careful threshold selection is crucial for maintaining an acceptable false positive rate, as traversing long texts is prone to accumulating false positives.
@@
 FPR FNR F1 IoU FPR FNR F1 IoU
-Llama-2-7b Full-text Detection 0.017 1.000 0.000 0.001 0.017 1.000 0.000 0.001
-WinMax 0.033 0.257 0.837 0.671 0.010 0.303 0.816 0.629
-FLSW-100 0.017 0.480 0.677 0.434 0.000 0.473 0.690 0.430
-FLSW-200 0.010 0.413 0.734 0.452 0.003 0.543 0.626 0.371
-FLSW-300 0.010 0.563 0.604 0.359 0.003 0.787 0.351 0.294
-FLSW-400 0.007 0.807 0.322 0.272 0.007 0.893 0.192 0.222
-WaterSeeker(Ours) 0.030 0.270 0.830 0.653 0.007 0.330 0.799 0.602
-Mistral-7b Full-text Detection 0.030 1.000 0.000 0.001 0.017 1.000 0.000 0.002
-WinMax 0.060 0.147 0.892 0.745 0.043 0.453 0.688 0.488
-FLSW-100 0.010 0.397 0.748 0.509 0.013 0.577 0.589 0.332
-FLSW-200 0.013 0.343 0.786 0.510 0.017 0.657 0.505 0.295
-FLSW-300 0.023 0.573 0.589 0.389 0.017 0.807 0.320 0.233
-FLSW-400 0.027 0.840 0.271 0.289 0.017 0.933 0.123 0.171
-WaterSeeker(Ours) 0.047 0.173 0.883 0.709 0.037 0.483 0.665 0.457
+Llama-2-7b Full-text Detection 0.000 1.000 [0.987] 0.000 [0.026] 0.000 0.000 1.000 [0.967] 0.000 [0.065] 0.001
+WinMax 0.017 0.273 0.834 0.661 0.017 0.410 0.734 0.537
+FLSW-100 0.003 0.473 0.688 0.448 0.003 0.547 0.622 0.363
+FLSW-200 0.003 0.440 0.716 0.417 0.000 0.577 0.595 0.338
+FLSW-300 0.007 0.683 0.479 0.313 0.000 0.783 0.356 0.265
+FLSW-400 0.003 0.897 0.187 0.230 0.003 0.933 0.125 0.191
+WaterSeeker(Ours) 0.017 0.303 0.813 0.629 0.010 0.433 0.719 0.510
+Mistral-7b Full-text Detection 0.000 1.000 [0.990] 0.000 [0.020] 0.000 0.000 1.000 [0.973] 0.000 [0.052] 0.001
+WinMax 0.010 0.277 0.835 0.656 0.013 0.370 0.767 0.562
+FLSW-100 0.000 0.500 0.667 0.433 0.000 0.497 0.670 0.411
+FLSW-200 0.000 0.410 0.742 0.452 0.003 0.557 0.613 0.360
+FLSW-300 0.003 0.573 0.597 0.342 0.003 0.827 0.295 0.254
+FLSW-400 0.003 0.810 0.318 0.244 0.007 0.943 0.107 0.195
+WaterSeeker(Ours) 0.007 0.287 0.829 0.642 0.010 0.390 0.753 0.542
 In Table 1, we present the detection performance of WaterSeeker alongside various baseline algorithms, evaluated on four datasets.
+Given that full-text detection lacks localization capabilities, its FNR is necessarily 1 under the evaluation rule requiring IoU $>\theta$ for successful detection.
+Therefore, we additionally report the FNR and corresponding F1 score for Full-text Detection without the IoU $>\theta$ constraint, highlighted in red.
 The results demonstrate that WaterSeeker achieves lower FNR and higher F1 score, significantly surpassing other watermark detection algorithms, including Full-text Detection, FLSW-100, FLSW-200, FLSW-300, and FLSW-400, while being comparable to WinMax.
@@
 However, subsequent experiments show that WinMax has extremely high time complexity (Section 6.4).
-Furthermore, the IoU results in the table reveal that WaterSeeker possesses a significantly higher localization capability compared to the baselines, again comparable to WinMax.
+Furthermore, the IoU [^1] For samples with no detected segments, the IoU is 0.
+For samples with multiple detected segments, take the max(IoU).
+results in the table reveal that WaterSeeker possesses a significantly higher localization capability compared to the baselines, again comparable to WinMax.
 6.3 Adaptability Analysis
@@
 Sections 4.2 and 5.4 analyze the theoretical time complexity of WinMax and WaterSeeker, respectively.
-This section will further delve into a more in-depth analysis based on experimental results.
+This section aims to offer a more comprehensive analysis based on experimental results, taking into account the constant factors preceding the complexity terms.
+Additionally, it will investigate the reasons behind the significant disparity in execution time of WinMax between KGW and Aar.
 In this experimental setting, let the time required to compute a score in the score list be $t_{1}$, and the time required to compute the watermark detection statistic within a local window be $t_{2}$.
@@
 FPR FNR F1 IoU FPR FNR F1 IoU
-Llama-2-7b w. local traverse 0.030 0.270 0.830 0.653 0.007 0.330 0.799 0.602
-w/o local traverse 0.010 0.333 0.795 0.565 0.007 0.400 0.747 0.509
-Mistral-7b w. local traverse 0.047 0.173 0.883 0.709 0.037 0.483 0.665 0.457
-w/o local traverse 0.030 0.267 0.832 0.619 0.027 0.510 0.646 0.400
+Llama-2-7b w. local traverse 0.017 0.303 0.813 0.629 0.010 0.433 0.719 0.510
+w/o local traverse 0.003 0.390 0.756 0.520 0.000 0.450 0.710 0.462
+Mistral-7b w. local traverse 0.007 0.287 0.829 0.642 0.010 0.390 0.753 0.542
+w/o local traverse 0.003 0.380 0.763 0.534 0.003 0.413 0.738 0.490
 7 Further Application
@@
 Appendix B Detailed Proof for Aar
-For the Aar method, we aim to analyze how the expected p-value for watermarked text varies with changes in the detection window size $W$.
+B.1 Expected p-value of watermarked text
+In this sub-section, we aim to analyze how the expected p-value for watermarked text varies with changes in the detection window size $W$.
 Recall the p-value calculation formula:
@@
 While this analysis does not strictly prove that the expectation of p-value is minimized exactly at $W=L$, it strongly suggests that the expectation of p-value is likely to reach its minimum value at or very near $W=L$ [^3] For a more rigorous proof, a deeper mathematical analysis of the GammaTransform function or numerical simulations would be necessary.
+B.2 Calculation of p-threshold to control false positive rate
 Similar to the analysis for KGW method, we also need to analyze the constraints on the p-value threshold $p^{*}$ when the false positive rate within the specified window is set to be lower than a target value $\alpha$.
-In the Aar method, the test statistic $S$ follows a Gamma distribution with shape parameter $W$ and scale parameter 1.
-For non-watermarked text, we have: $S\sim\text{Gamma}(W,1)$.
+For non-watermarked text, $u_{i}\sim\text{Uniform}([0,1])$.
+Consequently, the test statistic $S$ follows a Gamma distribution: $S\sim\text{Gamma}(W,1)$, where $W$ is the shape parameter and 1 is the scale parameter.
 The p-value is calculated using the Gamma CDF:
@@
 Solving this equation for $p^{*}$, we get $p^{*}=\alpha$, which is also a constant value for different $W$.
-Appendix C Detail of Threshold Selection in Full-text Detection
+Appendix C Detail of Threshold Selection Within the Specified Window
 A key role of threshold selection is to control the false positive rate.
@@
 Therefore, managing the false positive rate within the detection window is crucial in this scenario.
-Assuming the false positive rate of event $X$ is $\alpha$, let event $Y$ represent ”event $X$ has at least one false positive in 10,000 independent trials.
-” We know that $P(Y)=1-(1-\alpha)^{10000}$.
-If we want to control $P(Y)$ to be around 0.01, then $\alpha$ should be set to approximately $10^{-6}$.
-Next, we will explain how to set thresholds for both KGW and Aar to ensure the theoretical false positive rate in window detection is below $10^{-6}$.
+In the experiment, we set the targeted false positive rate $\alpha$ within the detection window to $10^{-6}$.
+C.1 Rationale for Setting $\alpha$ to $10^{-6}$
+Table 5: Simulated FPR of WaterSeeker using 10,000 samples for each watermarking method. The targeted false positive rate within the detection window is set to $10^{-6}$.
+Watermarking Method Simulated FPR
+KGW 0.0054
+Aar 0.0042
+WaterSeeker, WinMax, and FLSW all involve employing sliding windows for text traversal and conduct full-text detection within each window.
+As these windows overlap, they cannot be treated as independent, making it challenging to derive a theoretical upper bound for the document-level FPR from the targeted FPR within each window.
+Given this, we utilize large-scale data simulation to demonstrate that, with a targeted false positive rate of $10^{-6}$ within each window, our proposed method WaterSeeker maintains an acceptable false positive rate.
+For the KGW method, we set $\gamma=0.5$ in our experiments, meaning each token in non-watermarked text has a 0.5 probability of being green and 0.5 probability of being red.
+In the simulation, we generate 10,000 samples, each containing 10,000 tokens, with each token having a 0.5 probability of being 1 and 0.5 probability of being 0.
+For the Aar method, each token in non-watermarked text corresponds to $u_{i}\sim\text{Uniform}[0,1]$.
+In the simulation, we again generate 10,000 samples, each containing 10,000 tokens, with each token randomly assigned a floating-point number from $[0,1]$.
+We then apply WaterSeeker to detect watermarked segments within these samples, setting the targeted false positive rate within the detection window to $10^{-6}$.
+The large-scale simulation results in Table 5 demonstrate that WaterSeeker maintains a false positive rate of approximately 0.005, which is considered acceptable.
+For scenarios requiring more stringent FPR control, the targeted false positive rate can be adjusted downward.
+However, this inevitably compromises the detection rate, highlighting a key challenge in watermarked segment detection within large documents.
+C.2 Setting the Threshold to Achieve a Target False Positive Rate $\alpha$
 For KGW, as analyzed in Section 5.1, when the window size is large, we can approximate using the Central Limit Theorem, resulting in $z^{*}=\Phi^{-1}(1-\alpha)$.

### v3

@@
-WaterSeeker: Efficient Detection of Watermarked Segments in Large Documents
+WaterSeeker: Pioneering Efficient Detection of Watermarked Segments in Large Documents
 Abstract
@@
 Our code is available at https://github.com/THU-BPM/WaterSeeker.
+WaterSeeker: Pioneering Efficient Detection of Watermarked Segments in Large Documents
+Leyi Pan1,     Aiwei Liu 1,     Yijian Lu1,     Zitian Gao2,     Yichen Di1,     Lijie Wen1thanks: Corresponding author,     Irwin King3,     Philip S.
+Yu4 1Tsinghua University    2The University of Sydney 3The Chinese University of Hong Kong    4University of Illinois at Chicago panly24@mails.tsinghua.edu.cn, liuaw20@mails.tsinghua.edu.cn, wenlj@tsinghua.edu.cn
 1 Introduction
-As the quality of text generated by large language models (LLMs) improves, these models address various practical challenges while also raising concerns such as misinformation (Liu et al., 2023b; Chen & Shu, 2024) and copyright infringement (Rillig et al., 2023).
-To tackle these issues, LLM watermarking technology has emerged, embedding specific information (watermarks) during text generation for accurate detection through specialized algorithms.
-Existing watermarking methods mainly focus on distinguishing fully watermarked text from non-watermarked content (Kirchenbauer et al., 2023; Zhao et al., 2024; Liu et al., 2024b; Lu et al., 2024; Aaronson & Kirchner, 2022; Kuditipudi et al., 2024).
+As large language models (LLMs) generate high-quality text, they address practical challenges but also raise concerns such as misinformation (Liu et al., 2023b; Chen and Shu, 2024) and copyright infringement (Rillig et al., 2023).
+LLM watermarking technology has emerged to tackle these issues by embedding specific information (watermarks) during text generation, allowing for accurate detection through specialized algorithms.
+Existing watermarking methods primarily focus on distinguishing fully watermarked text from non-watermarked content (Kirchenbauer et al., 2023; Zhao et al., 2024; Liu et al., 2024b; Lu et al., 2024; Aaronson and Kirchner, 2022; Kuditipudi et al., 2024).
 However, in real-world scenarios, LLMs often generate only brief segments within longer documents, making it essential to detect these watermarked sections among large amounts of non-watermarked text.
-Most previous algorithms relied on computing statistics across the entire document for detection, which we refer to as full-text detection methods.
-For instance, some approaches categorize vocabulary into green and red lists, applying a bias towards the selection of green tokens during generation and using metrics like the z-score to measure their proportion throughout the entire text (Kirchenbauer et al., 2023; Zhao et al., 2024; Liu et al., 2024b).
-Other methods utilize pseudo-random numbers to guide token sampling, calculating correlations across the entire document during detection (Aaronson & Kirchner, 2022; Christ et al., 2024; Kuditipudi et al., 2024).
-Ultimately, the presence of a watermark is determined by whether these statistics exceed a specified threshold.
-However, these full-text detection methods struggle when small amounts of watermarked text are mixed with large volumes of natural text due to dilution effects, as illustrated in Figure 1(a).
+Most previous algorithms relied on computing statistics across the entire document for detection, which we refer to as full-text detection methods (detailed in Section 2).
+However, these full-text detection methods struggle when small watermarked sections are mixed with large volumes of natural text due to dilution effects, as illustrated in Figure 1(a).
 To the best of our knowledge, the only existing work addressing this scenario is the WinMax (Kirchenbauer et al., 2024) algorithm, which examines all possible window sizes and traverses the full text for each size.
 However, this method suffers from high time complexity.
-Additionally, our self-constructed baseline, the Fixed-Length Sliding Window (FLSW) method, faces the issue of watermark strength dilution due to inflexible window sizes, which negatively impacts detection performance.
+Additionally, our self-constructed baseline, the Fixed-Length Sliding Window (FLSW) method, faces the issue of watermark strength dilution due to inflexible window sizes, which lowers detection effectiveness.
 (a) Full-text detection methods often struggle in watermarked segment detection scenarios due to the dilution effect.
-(b) An illustration of the WaterSeeker algorithm, using “first locate, then detect” strategy to balance time complexity and detection performance.
+(b) An illustration of the WaterSeeker algorithm, using “first locate, then detect" strategy to balance time complexity and detection performance.
 Figure 1: An overview of watermark detection scenarios and the proposed WaterSeeker algorithm.
-To address these issues, we propose a novel watermark detection method named WaterSeeker, which is adaptable to various watermarking techniques.
-This method achieves a superior balance between time complexity and detection performance when identifying watermarked segments within large documents.
-The core concept of WaterSeeker is “first locate, then detect”.
-As illustrated in Figure 1(b), it initially employs a low-complexity anomaly points extraction algorithm to identify suspected watermark regions.
-This process narrows the detection target from a long text down to a small segment that encompasses the ground truth segment.
-Next, a local traversal is performed on the localization result.
-Within each window, full-text watermark detection is conducted, and the highest confidence result is compared to a threshold for the final determination.
-Theoretical analysis indicates that this coarse-to-fine process has the potential to achieve optimal detection performance while maintaining the lowest possible complexity for solving this problem.
-In the experiment, we compared the effectiveness of WaterSeeker and the baselines in detecting watermarked segments of random lengths inserted into large documents, as well as their time complexity.
-The results show that WaterSeeker significantly surpasses the baseline methods in effectively balancing time complexity and detection performance.
-Additionally, more detailed analysis includes an adaptability assessment of WaterSeeker under varying watermark strengths and segment lengths, as well as an evaluation of its two-stage utility via ablation study.
-Furthermore, we discuss WaterSeeker’s contribution to building more interpretable AI detection systems.
+To address these issues, we propose a novel watermark detection method called WaterSeeker.
+WaterSeeker follows a "first locate, then detect" approach, as shown in Figure 1(b).
+It initially employs a low-complexity anomaly points extraction algorithm to identify suspected watermark regions, narrowing the detection target from a long text to a small segment encompassing the ground truth segment.
+Next, a local traversal is performed on the localization result, conducting full-text watermark detection within each window and comparing the highest confidence result to a threshold for the final determination.
+Theoretical analysis suggests that this coarse-to-fine process has the potential to achieve optimal detection performance while maintaining the lowest possible complexity for solving this problem.
+In the experiment, we compared the effectiveness and time complexity of WaterSeeker with baseline methods for detecting watermarked segments in large documents.
+WaterSeeker significantly outperformed the baselines in balancing time complexity and detection performance.
+We also assessed its adaptability to different watermark strengths and segment lengths, and evaluated its two-stage utility through an ablation study.
 In summary, the contributions of this work are as follows:
-  • We comprehensively define the new scenario of detecting watermarked segments in large documents, specifying the inputs and outputs of the algorithms designed to address this problem, as well as the evaluation metrics and the construction process for the test dataset.
-  • We propose WaterSeeker, a novel and general watermark detection method that effectively identifies watermarked segments in large documents, addressing the challenges posed by dilution effects in non-watermarked text.
-  • WaterSeeker achieves a superior balance between time complexity and detection effectiveness, with strong adaptability to varying watermark strengths and lengths, significantly outperforming existing baselines.
-  • WaterSeeker not only provides accurate detection results but also offers precise localization outcomes, contributing to the development of more interpretable AI detection systems.
+  • We comprehensively define a new scenario: detecting watermarked segments in large documents. This includes specifying algorithm inputs/outputs, evaluation metrics, and how to create test datasets.
+  • We propose WaterSeeker, a general watermark detection method that effectively identifies watermarked segments in large documents, tackling the issues caused by dilution effects.
+  • WaterSeeker outperforms baselines in achieving a superior balance between time complexity and detection effectiveness.
+  • WaterSeeker not only provides accurate detection results but also pinpoints watermark locations precisely, helping to create more explainable AI detection systems (detailed in Appendix H).
 2 Related Work
-Watermarking for large language models (LLMs) embeds subtle, algorithmically detectable signals in outputs to identify LLM-generated text, making it crucial for mitigating potential misuse of LLMs.
-Currently, mainstream watermarking methods involve intervening in the inference phase of LLMs by modifying the logits output or influencing the token sampling process to produce watermarked text (Liu et al., 2023a; Pan et al., 2024; Liu et al., 2024c).
-The KGW method (Kirchenbauer et al., 2023) modifies logits by categorizing the vocabulary into green and red lists, applying a bias towards green tokens during text generation.
+Currently, mainstream LLM watermarking methods involve modifying the inference phase by altering logits or influencing token sampling (Liu et al., 2023a; Pan et al., 2024; Liu et al., 2024c).
+The KGW family (Kirchenbauer et al., 2023; Zhao et al., 2024; Hu et al., 2024; Liu et al., 2024b; He et al., 2024; Lu et al., 2024; Liu et al., 2024a) categorizes vocabulary into green and red lists, biasing towards green tokens during generation.
 The bias value is typically determined by the parameter $\delta$, which reflects the watermark strength.
-For detection, it first labels each token as red or green and calculates the z-score of the entire text to reflect the proportion of green tokens; if this proportion exceeds a specified threshold, the text is classified as watermarked.
-In subsequent work, numerous enhancements have been suggested to improve text quality (Hu et al., 2024; Wu et al., 2023; Takezawa et al., 2023), increase information capacity (Wang et al., 2024; Yoo et al., 2023; Fernandez et al., 2023), boost robustness against watermark removal attacks (Zhao et al., 2024; Liu et al., 2024b; Ren et al., 2024; He et al., 2024), adapt to low-entropy conditions (Lee et al., 2023; Lu et al., 2024), and facilitate public detection (Liu et al., 2024a; Fairoze et al., 2023).
-On the other hand, the Aar method (Aaronson & Kirchner, 2022) proposes utilizing pseudo-random sequences to guide the token sampling process.
-During each token sampling step, a pseudo-random vector $u\sim\text{Uniform}([0,1])^{|V|}$ is generated based on the previously sampled tokens.
-Assuming that the probability vector provided by the LLM is $p$, the method performs exponential sampling to select the token $i$ that maximizes $u_{i}^{1/p_{i}}$.
-Here, the watermark strength is reflected by a sampling temperature, which affects the strictness of the exponential sampling.
+Detection involves calculating the z-score of green tokens among the entire document; exceeding a threshold indicates watermarking.
+On the other hand, the Aar family (Aaronson and Kirchner, 2022; Christ et al., 2024; Kuditipudi et al., 2024) uses pseudo-random sequences to guide token sampling.
+It generates a pseudo-random vector $u\sim\text{Uniform}([0,1])^{|V|}$ based on previous tokens and selects the token $i$ maximizing $u_{i}^{1/p_{i}}$, where $p$ is the LLM’s probability vector.
+Watermark strength is controlled by sampling temperature.
 For detection, it also employs global statistics: it sums the correlation values of each token with the pseudo-random vector and then performs gamma transformation to derive the detection confidence.
-Building on this, Kuditipudi et al.
-(2024) introduced the use of edit distance to calculate correlation, while Hou et al.
-(2024a) and Hou et al.
-(2024b) incorporated contextual semantic information to guide the sampling process, thereby enhancing detection robustness.
+Details of the KGW and Aar watermarking algorithms can be found in Appendix A.
 Despite the high accuracy of watermarking algorithms for detecting LLM-generated text, previous studies have primarily focused on distinguishing between fully watermarked and non-watermarked text, overlooking the possibility that LLMs may only generate small segments within large documents.
 In such cases, watermark detection algorithms based on full-text statistics can fail due to the dilution effect.
-A few studies have mentioned copy-paste attack (Kirchenbauer et al., 2024; Yoo et al., 2023; Wang et al., 2024), where a certain proportion of watermarked text is mixed into non-watermarked content, resembling our scenario.
+A few studies have mentioned copy-paste attack (Kirchenbauer et al., 2024; Yoo et al., 2023; Wang et al., 2024), which involves mixing a portion of watermarked text with non-watermarked content, similar to our scenario.
 Yoo et al.
 (2023) and Wang et al.
-(2024) tested the robustness of their methods against copy-paste attacks by mixing 10% to 50% watermarked text into non-watermarked text.
-However, since they did not specifically design detection mechanisms for this scenario, their results indicated a lack of robustness against such attack.
+(2024) evaluated their methods’ robustness against copy-paste attacks by combining 10% to 50% watermarked text with non-watermarked text.
+However, as they did not develop specific detection mechanisms for this situation, their findings showed that their methods were not robust against this type of attack.
 In existing studies, WinMax (Kirchenbauer et al., 2024) is the only watermark detection algorithm specifically designed for this scenario.
@@
 While this method shows some effectiveness, its high time complexity renders it impractical for real-world applications.
-Therefore, to achieve effective detection of watermarked segments inserted into large documents while maintaining controllable time complexity, we propose WaterSeeker, a novel and general method that uses “first locate then detect” strategy.
-We validate its superiority through comparisons with full-text detection, WinMax, and a self-constructed fixed-length sliding window detection method.
+Therefore, to achieve effective detection of watermarked segments inserted into large documents while maintaining controllable time complexity, we propose WaterSeeker, a novel and general method that uses “first locate then detect" strategy.
 3 Problem Formulation
 3.1 Definition
-The problem of detecting watermarked segments in a long document is defined as follows: Let $N$ be a long text with a randomly inserted watermarked segment of length $L$ (where $L$ is a random variable).
+The problem of detecting watermarked segments in a long document is defined as follows: Let $N$ be a long text with a randomly inserted watermarked segment of length $L$.
 We denote the starting index of the watermarked segment in $N$ as $s$ and the ending index as $e$, such that: $L=e-s+1$.
@@
 […]
+[…]
 Here, ‘has_watermark’ is a boolean value indicating the presence of a watermark, and ‘indices’ is a list of pairs $(s^{\prime},e^{\prime})$ representing the start and end indices of detected watermarked segments.
 3.2 Evaluation
-A watermark is considered successfully detected if:
 […]
-  • There exists at least one detected segment $(s^{\prime},e^{\prime})$ such that the Intersection over Union (IoU) with the ground truth segment $(s,e)$ exceeds a specified threshold $\theta$:
-    $\text{IoU}=\cfrac{L_{\text{intersection}}}{L_{\text{union}}}>\theta.$ (1)
-Based on this evaluation, we will report the following metrics of the binary classification result: False Positive Rate (FPR), False Negative Rate (FNR), and F1 Score.
-Additionally, we will include the average IoU between the detected segments and ground truth segments as a supplementary metric, further demonstrating the accuracy of the localization.
+$\text{IoU}=\cfrac{L_{\text{intersection}}}{L_{\text{union}}}>\theta.$
+Based on this criteria, we will report the following metrics of the binary classification result: False Positive Rate (FPR), False Negative Rate (FNR), and F1 Score.
+Moreover, we will include the average IoU between the detected segments and ground truth segments as a supplementary metric, further demonstrating the accuracy of the localization.
 4 Baseline Methods
@@
 4.2 WinMax
-WinMax (Kirchenbauer et al., 2024) is a watermark detection algorithm proposed on the basis of KGW to address copy-paste attack.
-It involves iterating through all possible window sizes, and for each window size, the entire text is traversed to calculate the z-score for each local window, taking the maximum z-score and comparing it against a specified threshold.
+WinMax (Kirchenbauer et al., 2024) involves iterating through all possible window sizes, and for each window size, the entire text is traversed to calculate the z-score for each local window, taking the maximum z-score and comparing it against a specified threshold.
 In practical applications, a minimum window size $W_{min}$ and a maximum window size $W_{max}$ are typically set to control the number of iterations during the traversal.
 The detection process can be described by the following formula:
-$z_{\text{win-max}}=\max_{w\in[W_{min},W_{max}]}\left(\max_{i}\left(\frac{|s|_{G,i}-\gamma w}{\sqrt{\gamma(1-\gamma)w}}\right)\right),$ (2)
+$z_{\text{win-max}}=\max_{w\in[W_{min},W_{max}]}\left(\max_{i}\left(\frac{|s|_{G,i}-\gamma w}{\sqrt{\gamma(1-\gamma)w}}\right)\right),$
 where $w$ is the length of the local window, and $|s|_{G,i}$ represents the count of green tokens within the $i$-th local window of length $w$.
-For other watermarking algorithms, such as Aar, a similar approach can also be implemented.
-The time complexity of the WinMax detection method is relatively high, costing $O((W_{max}-W_{min})\times N)$, where $N$ is the total length of the text.
+The time complexity of WinMax is $O((W_{max}-W_{min})\times N)$, where $N$ is the total length of the text.
 Due to the uncertain length of the watermarked segment inserted into the document, if there is a large difference between $W_{max}$ and $W_{min}$, the worst-case complexity can reach $O(N^{2})$.
 4.3 Fix-Length Sliding Window
-Fix-Length Sliding Window (FLSW) detection method is a self-constructed, straightforward approach to address the watermark detection problem in this scenario.
-It involves using a fixed-length window to traverse the entire text, calculating statistics within the local window.
-If the statistics within a local window exceed a specified threshold, the variable ‘has_watermarked’ is set to true.
-For localization, during the traversal, each pair of indices $(s,e)$ that meets the condition is recorded.
-Finally, a contiguous method with a certain tolerance is employed to provide the localization results.
-The pseudocode for the aforementioned baselines can be found in Appendix A.
-Although the concept of FLSW is relatively simple and straightforward, the fixed-length nature makes it unable to adapt to varying watermark lengths.
-This limitation can result in dilution effects when the window size is either too large or too small.
-In Section 5.1, we will analyze this dilution effect in detail using KGW (Kirchenbauer et al., 2023) and Aar (Aaronson & Kirchner, 2022) as examples, providing theoretical support for WaterSeeker.
+Fix-Length Sliding Window (FLSW) is a self-constructed method that uses a fixed-length window to traverse the text, calculating statistics within each local window.
+The text is flagged as watermarked if any statistic score exceeds the threshold.
+For localization, the method records each pair of indices $(s,e)$ meeting the condition and uses a contiguous segment merging method with tolerance.
+While straightforward, FLSW struggles with varying watermark lengths due to its fixed-length nature, leading to dilution effects.
+Section 5.1 analyzes this effect using KGW (Kirchenbauer et al., 2023) and Aar (Aaronson and Kirchner, 2022) as examples, providing theoretical support for WaterSeeker.
 5 Proposed Method: WaterSeeker
-5.1 Theoretical Foundation: Gold Index is the Best
-In this section, we will conduct a theoretical analysis to demonstrate that using the actual start and end indices (the gold index) for watermark detection yields the highest expected detection rate, using KGW and Aar as examples.
-As mentioned in Section 2, most watermarking algorithms are slight modifications of these two methods, maintaining the fundamental principles of statistical calculation.
-Therefore, analyzing these two approaches provides a broad applicability.
-For the KGW method, assuming $\gamma_{1}>\gamma$ is the proportion of green tokens in the watermarked part. Let’s analyze the effect of window size $W$ on this statistic:
-  • When $W<L$ (window size smaller than watermark length), we have $E[|s|_{G}]=W\gamma_{1}$. Therefore,
-    $E[z_{W}]=\frac{W\gamma_{1}-\gamma W}{\sqrt{\gamma(1-\gamma)W}}=\sqrt{W}\cdot\frac{\gamma_{1}-\gamma}{\sqrt{\gamma(1-\gamma)}},$ (3)
-    where the expectation of z-score is proportional to $\sqrt{W}$.
-  • When $W>L$ (window size larger than watermark length), we have $E[|s|_{G}]=L\gamma_{1}+(W-L)\gamma$. Therefore
-    $E[z_{W}]=\frac{L\gamma_{1}+(W-L)\gamma-\gamma W}{\sqrt{\gamma(1-\gamma)W}}=\frac{L(\gamma_{1}-\gamma)}{\sqrt{\gamma(1-\gamma)W}},$ (4)
-    where the expectation of z-score is proportional to $1/\sqrt{W}$, decreasing as $W$ increases.
+5.1 Theoretical Basis: Gold Index is the Best
+This section provides the theoretical foundation of WaterSeeker, showing that using actual start and end indices (gold index) for watermark detection achieves the highest expected detection rate.
+We use KGW and Aar as examples, since most watermarking algorithms are slight modifications of these two methods.
+Thus, analyzing these approaches offers broad applicability.
+For the KGW method, assuming $\gamma_{1}>\gamma$ is the proportion of green tokens in the watermarked part.
+Let’s analyze the effect of window size $W$ on this statistic: (1) When $W<L$ (window size smaller than watermark length):
+$E[z_{W}]=\frac{W\gamma_{1}-\gamma W}{\sqrt{\gamma(1-\gamma)W}}=\sqrt{W}\cdot\frac{\gamma_{1}-\gamma}{\sqrt{\gamma(1-\gamma)}}.$
+(2) When $W>L$ (window size larger than watermark length):
+$E[z_{W}]=\frac{L\gamma_{1}+(W-L)\gamma-\gamma W}{\sqrt{\gamma(1-\gamma)W}}=\frac{L(\gamma_{1}-\gamma)}{\sqrt{\gamma(1-\gamma)W}}.$
 From this, we can conclude that when $W=L$, the z-score reaches its maximum.
@@
 We start with the binomial distribution $B(W,\gamma)$ that describes the number of green tokens in a window of size $W$.
-For large sample sizes (typically $W>50$), we can apply the Central Limit Theorem to approximate this binomial distribution with a normal distribution $N(W\gamma,W\gamma(1-\gamma))$.
-We then standardize this to a z-score, which follows the standard normal distribution $N(0,1)$.
-The false positive rate alpha represents the area in the right tail of this distribution beyond $z^{*}$.
-Mathematically, this is expressed as $\alpha=1-\Phi(z^{*})$, where $\Phi$ is the cumulative distribution function (CDF) of the standard normal distribution.
-Solving for $z^{*}$, we get:
-$z^{*}=\Phi^{(-1)}(1-\alpha),$ (5)
-which is a constant value for different $W$.
+For large sample sizes (typically $W>50$), we can approximate this as $|s|_{G}\sim N(W\gamma,W\gamma(1-\gamma))$.
+Therefore, we have $z\sim N(0,1)$.
+The false positive rate $\alpha$ represents the area in the right tail of this distribution beyond $z^{*}$.
+Mathematically, this is expressed as $\alpha=1-\Phi(z^{*})$.
+Solving for $z^{*}$, we get: $z^{*}=\Phi^{(-1)}(1-\alpha)$, which is a constant value for different $W$.
 For the Aar method, the p-value is calculated by applying a gamma transformation to the sum of correlation values of all the tokens:
 […]
-Due to the complexity of the Gamma-Transformation, the detailed proof can be found in Appendix B.
+We denote $\mu=E[\log(\frac{1}{1-u_{i}})]$ as the expected correction value for a single token.
+Similar to KGW, we assume $\mu_{1}>\mu$ represents the expected correlation value for the watermarked portion.
+Due to the complexity of the Gamma-Transformation, the detailed proof can be found in Appendix C.
 Based on the above proof, we simulated the expected score of the statistic and the corresponding threshold as they vary with $W$ for $L=200$ using real data, as shown in Figure 2.
 It demonstrates that when the window size corresponds to the gold index, the highest detection rate can be achieved while maintaining an acceptable FPR.
-Therefore, the watermark dilution effect refers to the situation where the window used for detection is not optimal, resulting in the computed statistics representing watermark strength not reaching their peak or minimum, which in turn affects detection performance.
+Therefore, the watermark dilution effect refers to the situation where the window used for detection is not optimal, resulting in the computed statistics not reaching their peak or minimum, which in turn affects detection performance.
 (a) KGW, $\alpha=10^{-6},\gamma=0.5,\gamma_{1}=0.75$
 (b) Aar, $\alpha=10^{-6},\mu=1.0,\mu_{1}=1.6$
-Figure 2: Expected score of the statistics and the corresponding threshold across varying $W$.
+Figure 2: Expected score of the statistics and the corresponding threshold across various $W$.
 5.2 Suspicious Region Localization
@@
 Below is the detailed process of the localization algorithm:
-  1. 1.
-    Score List Computation: We utilize a small sliding window (i.e., $W=50$) to efficiently traverse the entire text.
-For each position of the window, we compute a statistical measure that reflects the average watermark intensity.
-For example, in the KGW method, this statistic is represented as $|s|_{G}/W$, indicating the density of green tokens inside the window.
-Meanwhile, in the Aar method, it is calculated as $\cfrac{\sum\log(1/(1-u))}{W}$, representing the average correlation value between tokens and the pseudo-random sequence within the window.
-This results in a score list $s$ of length $N-W+1$, where $s_{i}$ represents the average watermark intensity of $n_{i}$ to $n_{i}+W$.
-  2. 2.
-    Anomaly Extraction: In this step, we design an anomaly extraction algorithm to identify outliers from the score list.
+(1) Score List Computation: A small sliding window (i.e., $W=50$) is used to traverse the entire text.
+For each window position, a statistical measure is computed to reflect the average watermark intensity.
+In the KGW method, this statistic is $|s|_{G}/W$, representing the density of green tokens within the window.
+In the Aar method, it is calculated as $\frac{\sum\log(1/(1-u))}{W}$, representing the average correlation value between tokens and the pseudo-random sequence within the window.
+This results in a score list $s$ of length $N-W+1$, where $s_{i}$ represents the average watermark intensity from $n_{i}$ to $n_{i}+W$.
+(2) Anomaly Extraction: In this step, we design an anomaly extraction algorithm to identify outliers from the score list.
 We calculate the mean of the score list, the top-k mean, and the standard deviation, denoted as $s_{\text{mean}}$, $s_{\text{top-k-mean}}$, and $s_{\text{std}}$, respectively.
 The extracted outliers are the points that satisfy the following condition, where $\theta_{1}$ and $\theta_{2}$ are hyperparameters:
-    $\text{score}>s_{\text{mean}}+\max((s_{\text{top-k-mean}}-s_{\text{mean}})\cdot\theta_{1},s_{\text{std}}\cdot\theta_{2}).$ (7)
+$\text{score}>s_{\text{mean}}+\max((s_{\text{top-k-mean}}-s_{\text{mean}})\cdot\theta_{1},s_{\text{std}}\cdot\theta_{2}).$
 In samples containing a watermarked segment, the first term in the max expression becomes significant.
@@
 In this case, to reduce false positives, a point must deviate from the mean by at least $s_{\text{std}}\cdot\theta_{2}$ to be considered an anomaly.
-  3. 3.
-    Fragment Connection: In this step, we use a connecting method with a certain tolerance to link nearby outliers, then filter out segments that are too short, returning a list of indices.
-The scores within watermarked segments tend to be relatively stable and close to the top-k mean.
-Concurrently, the scores corresponding to segments outside the watermark also remain stable and approximately equal to the overall mean (because a large volume of natural text dilutes the overall watermark strength).
-As a result of these characteristics, when we use $(s_{\text{top-k-mean}}-s_{\text{mean}})\cdot\theta_{1}\quad(\theta_{1}<1)$ as a criterion for distinction, the extracted abnormal segment’s start and end points ($s^{\prime}$ and $e^{\prime}$) generally satisfy the following conditions: $s^{\prime}\in(s-W,s)$ and $e^{\prime}\in(e,e+W)$.
-This ensures that the extracted suspicious watermarked regions are likely to cover the actual segments while keeping the starting and ending deviations within a window size.
-Algorithm 1 WaterSeeker Algorithm
-1:procedure SuspiciousRegionLocalization(tokens, $W$, $k$, $\theta_{1}$, $\theta_{2}$)
-2:     Compute score list using sliding window
-3:     Calculate $s_{\text{mean}}$, $s_{\text{top-k-mean}}$, and $s_{\text{std}}$
-4:     Detect anomalies using threshold:
-5:     $s_{\text{mean}}+\max((s_{\text{top-k-mean}}-s_{\text{mean}})\cdot\theta_{1},s_{\text{std}}\cdot\theta_{2})$
-6:     Connect nearby outliers and filter short segments
-7:     return filteredSegments
-8:end procedure
-9:procedure LocalTraverseDetection(tokens, suspiciousRegions, $W$, threshold)
-10:     hasWatermark $\leftarrow$ False
-11:     indices $\leftarrow$ [ ]
-12:     for $(s^{\prime},e^{\prime})$ in suspiciousRegions do
-13:         maxScore = $\infty$- ∞, maxIndice = ()
-14:         for $s\in[s^{\prime},s^{\prime}+W)$ and $e\in(e^{\prime}-W,e^{\prime}]$ do
-15:              if DetectWatermark(tokens[$s:e$])$>$ maxScore then
-16:                  maxScore = DetectWatermark(tokens[$s:e$])
-17:                  maxIndice = $(s,e)$
-18:              end if
-19:         end for
-20:         if maxScore $>$ threshold then
-21:              hasWatermark $\leftarrow$ True
-22:              indices.append(maxIndice)
-23:         end if
-24:     end for
-25:     return hasWatermark, indices
-26:end procedure
-27:procedure WaterSeeker(tokens, $W$, $k$, $\theta_{1}$, $\theta_{2}$, threshold)
-28:     suspiciousRegions $\leftarrow$ SuspiciousRegionLocalization(tokens, $W$, $k$, $\theta_{1}$, $\theta_{2}$)
-29:     hasWatermark, indices $\leftarrow$ LocalTraverseDetection(tokens, suspiciousRegions, W, threshold)
-30:     return hasWatermark, indices
-31:end procedure
-5.3 Local Traverse Detection
-After obtaining the localization results from the first step, the second phase (fine step), local traverse detection, conducts a more detailed verification.
-For each $(s^{\prime},e^{\prime})$ pair in the localization results, traverse inwards through the segments where the start point falls within $[s^{\prime},s^{\prime}+W)$ and the end point within $(e^{\prime}-W,e^{\prime}]$.
-Based on the previous analysis of the ranges of $s^{\prime}$ and $e^{\prime}$, this traversal has a high probability of reaching the gold index.
-Perform full-text detection sequentially on these segments, and select the statistic with the most significance (e.g., for KGW, choose the segment with the largest z-score, and for Aar, the segment with the smallest p-value) to compare with the threshold.
-The pseudocode for the entire WaterSeeker algorithm can be found in Algorithm 1.
-5.4 Time Complexity Analysis
-Time Complexity Analysis of WaterSeeker:
-  1. 1.
-    Suspicious Region Localization: $O(N)$.
-This step involves three sub-steps: score list computation, anomaly detection, and fragment connection.
-Each of these sub-steps has a time complexity of $O(N)$, where $N$ is the length of the text.
-  2. 2.
-    Local Traverse Detection: $O(W^{2})$. For each suspicious region identified, this step examines $W^{2}$ windows, where $W$ is the window size.
-The total time complexity of WaterSeeker is $O(N+W^{2})$.
-In practice, $W^{2}$ is typically maintained at a value lower than $T$, as even a slightly larger window (i.e., $W=50$) is sufficient to provide a relatively smooth and low-noise representation of the surrounding watermark intensity.
-Therefore, the dominant term in most cases is $O(N)$, making the overall time complexity of WaterSeeker $O(N)$.
-Lower Bound Complexity Analysis for the Problem: To detect watermarked segments in a long text, any algorithm must examine each token in the text at least once.
-This necessitates at least one full pass through the text, establishing a lower bound of $\Omega(N)$ for the time complexity of this problem.
-In conclusion, the WaterSeeker algorithm, with its $O(N)$ complexity in typical cases, achieves a time complexity that matches the theoretical lower bound of the problem, indicating that it is asymptotically optimal for the task of detecting watermarked segments in long texts.
-6 Experiment
-6.1 Experiment Settings
-Watermarking Methods and Language Models: We selected two representative watermarking algorithms, KGW (Kirchenbauer et al., 2023) and Aar (Aaronson & Kirchner, 2022), each with three strength levels.
-For KGW, the strength is determined by parameter $\delta$, with values 2.0, 1.5, and 1.0 (from strong to weak).
-For Aar, the strength is controlled by parameter temperature, with values 0.3, 0.2, and 0.1 (from strong to weak).
-We used Llama-2-7b (Touvron et al., 2023) and Mistral-7b (Jiang et al., 2024) as generation models.
-Dataset Construction: Following previous studies, the C4 dataset (Raffel et al., 2020) was used for prompts, with the first 30 tokens of each entry serving as input.
-Watermarked segments of random length (100 to 400 tokens) were generated using randomly selected watermark strengths.
-For positive examples, one such segment was randomly inserted into each 10,000-token Wikipedia passage (Foundation, ).
-Negative examples consist of unmodified 10,000-token Wikipedia corpus.
-Based on this procedure, four datasets were created, each containing 300 positive and 300 negative examples: KGW-llama, KGW-mistral, Aar-llama, and Aar-mistral.
-These datasets employ different combinations of watermarking methods and LLMs.
-The specific distributions of watermark strengths and lengths are detailed in Appendix D.
-It is important to note that watermarked segments with higher strength and longer length are inherently easier to detect and demonstrate better robustness against the dilution effect.
-Therefore, to showcase the superiority and adaptability of WaterSeeker, each dataset incorporates watermarked segments of varying strengths and lengths.
-Baselines: As introduced in Section 4, Full-text Detection and WinMax (Kirchenbauer et al., 2024) are chosen, as well as Fix-Length Sliding Window method with $W$ of 100, 200, 300 and 400.
-Hyper-parameters: The parameters related to WaterSeeker are as follows: $W=50,k=20$, with a tolerance for fragment connection set to 100.
-The parameter $\theta_{1}$ is set to 0.5, and $\theta_{2}$ is set to 1.5 for both algorithms.
-The threshold selection within the specified window is detailed in Appendix C.
-Notably, careful threshold selection is crucial for maintaining an acceptable false positive rate, as traversing long texts is prone to accumulating false positives.
-Evaluation-related parameters: $\theta$ mentioned in Section 3.2 is set to 0.5.
-6.2 Detection Performance Analysis
-Table 1: We compared the detection performance of WaterSeeker with other methods, including Plain Detect, WinMax (Kirchenbauer et al., 2024), and FLSW.
-The reported metrics include False Positive Rate (FPR), False Negative Rate (FNR), F1 score, and the average Intersection over Union (IoU) between detected segments and ground truth segments in positive samples.
+(3) Fragment Connection: In this step, we use a connecting method with a certain tolerance to link nearby outliers, then filter out segments that are too short, returning a list of indices.
+Table 1: We compared the detection performance of WaterSeeker with other methods, including Full-text Detection, WinMax (Kirchenbauer et al., 2024), and FLSW.
+The reported metrics include FPR, FNR, F1 score, and the average IoU between detected segments and ground truth segments in positive samples.
 Model Method KGW Aar
-FPR FNR F1 IoU FPR FNR F1 IoU
+FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$ FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$
 Llama-2-7b Full-text Detection 0.000 1.000 [0.987] 0.000 [0.026] 0.000 0.000 1.000 [0.967] 0.000 [0.065] 0.001
@@
 FLSW-400 0.003 0.897 0.187 0.230 0.003 0.933 0.125 0.191
-WaterSeeker(Ours) 0.017 0.303 0.813 0.629 0.010 0.433 0.719 0.510
+WaterSeeker(Ours) 0.017 0.313 0.806 0.624 0.010 0.440 0.713 0.507
 Mistral-7b Full-text Detection 0.000 1.000 [0.990] 0.000 [0.020] 0.000 0.000 1.000 [0.973] 0.000 [0.052] 0.001
@@
 FLSW-100 0.000 0.500 0.667 0.433 0.000 0.497 0.670 0.411
-FLSW-200 0.000 0.410 0.742 0.452 0.003 0.557 0.613 0.360
-FLSW-300 0.003 0.573 0.597 0.342 0.003 0.827 0.295 0.254
+FLSW-200 0.000 0.410 0.742 0.452 0.003 0.560 0.610 0.359
+FLSW-300 0.003 0.577 0.593 0.341 0.003 0.827 0.295 0.254
 FLSW-400 0.003 0.810 0.318 0.244 0.007 0.943 0.107 0.195
-WaterSeeker(Ours) 0.007 0.287 0.829 0.642 0.010 0.390 0.753 0.542
-In Table 1, we present the detection performance of WaterSeeker alongside various baseline algorithms, evaluated on four datasets.
-Given that full-text detection lacks localization capabilities, its FNR is necessarily 1 under the evaluation rule requiring IoU $>\theta$ for successful detection.
-Therefore, we additionally report the FNR and corresponding F1 score for Full-text Detection without the IoU $>\theta$ constraint, highlighted in red.
-The results demonstrate that WaterSeeker achieves lower FNR and higher F1 score, significantly surpassing other watermark detection algorithms, including Full-text Detection, FLSW-100, FLSW-200, FLSW-300, and FLSW-400, while being comparable to WinMax.
-Based on the principle that WinMax is guaranteed to reach the gold index by evaluating all possible windows, it represents the upper bound of detection performance.
-However, subsequent experiments show that WinMax has extremely high time complexity (Section 6.4).
-Furthermore, the IoU [^1] For samples with no detected segments, the IoU is 0.
-For samples with multiple detected segments, take the max(IoU).
-results in the table reveal that WaterSeeker possesses a significantly higher localization capability compared to the baselines, again comparable to WinMax.
-6.3 Adaptability Analysis
-While the four datasets used in the main experiment incorporated a mix of watermarked segments with varying strengths and lengths, this section presents a more granular comparison of the detection capabilities of WaterSeeker and the FLSW method (FLSW-100, FLSW-200, FLSW-300, and FLSW-400) across specific watermark strengths and lengths.
-As shown in Figure 3, the performance of different detection algorithms exhibits a consistent trend: watermarked segments with higher strengths and longer lengths are more easily detected.
-Moreover, it is evident that FLSW performs well within a length range close to its window size, but shows poor adaptability in other ranges.
-This limitation arises from its fixed-length nature, which leads to dilution effects when detecting longer or shorter watermarked segments, thereby reducing detection performance.
-In contrast, WaterSeeker exhibits good adaptability across different length ranges by using anomaly extraction techniques.
-Additionally, it incorporates a top-k score mechanism to adapt to varying watermark strengths.
+WaterSeeker(Ours) 0.007 0.290 0.827 0.641 0.010 0.390 0.753 0.542
+The scores within watermarked segments are relatively stable and close to the top-k mean, while scores outside the watermark remain stable and close to the overall mean due to the dilution effect of large natural text.
+When using $(s_{\text{top-k-mean}}-s_{\text{mean}})\cdot\theta_{1}\quad(\theta_{1}<1)$ as a distinction criterion, the extracted abnormal segment’s start and end points ($s^{\prime}$ and $e^{\prime}$) generally satisfy $s^{\prime}\in(s-W,s)$ and $e^{\prime}\in(e,e+W)$.
+This ensures that the extracted suspicious watermarked regions likely cover the actual segments, with starting and ending deviations within a window size.
+5.3 Local Traverse Detection
+After obtaining the localization results from the first step, the second phase (fine step), local traverse detection, conducts a more detailed verification.
+For each $(s^{\prime},e^{\prime})$ pair in the localization results, traverse inwards through the segments where the start point falls within $[s^{\prime},s^{\prime}+W)$ and the end point within $(e^{\prime}-W,e^{\prime}]$.
+Based on the previous analysis of the ranges of $s^{\prime}$ and $e^{\prime}$, this traversal has a high probability of reaching the gold index.
+Perform full-text detection sequentially on these segments, and select the statistic with the most significance to compare with the threshold.
+The pseudocode for the entire WaterSeeker algorithm can be found in Algorithm 3.
+5.4 Time Complexity Analysis
+Time Complexity of WaterSeeker: Suspicious Region Localization has a time complexity of $O(N)$.
+This step involves score list computation, anomaly detection, and fragment connection, each with a time complexity of $O(N)$, where $N$ is the text length.
+Local Traverse Detection has a time complexity of $O(W^{2})$.
+For each suspicious region, this step examines $W^{2}$ windows, where $W$ is the window size.
+The total time complexity of WaterSeeker is $O(N+W^{2})$.
+In practice, $W^{2}$ is typically kept lower than $N$, as a slightly larger window (i.e., $W=50$, detailed in Appendix E) suffices for a smooth and low-noise representation of the surrounding watermark intensity.
+Thus, the overall time complexity of WaterSeeker is $O(N)$.
+Lower Bound Complexity for the Problem: To detect watermarked segments in a long text, any algorithm must examine each token in the text at least once.
+This requirement establishes a lower bound of $\Omega(N)$ for the time complexity of the problem, as at least one full pass through the text is necessary.
+Consequently, the WaterSeeker algorithm achieves a time complexity that matches the theoretical lower bound of the problem.
+6 Experiment
 (a) KGW, $\delta=2.0$
@@
 Figure 3: This figure compares the detection performance of WaterSeeker and FLSW under varying watermark lengths and strengths, using Llama-2-7b as the generation model.
-6.4 Time Complexity Analysis
-Table 2: Average detection time per sample for various detection methods ($T\approx 10,000$, Unit: s).
+Table 2: Average detection time per sample for various detection methods ($N\approx 10,000$, Unit: s).
 Full-text Detection WinMax FLSW-100 FLSW-200 FLSW-300 FLSW-400 WaterSeeker
@@
 Aar 0.54 2733.78 1.62 1.65 1.66 1.64 1.68
-Table 2 lists the average time spent per sample for WaterSeeker and other baseline algorithms during detection.
-The results show that the time taken by WaterSeeker is comparable to that of full-text detection and FLSW, and significantly lower than the time required by WinMax.
-Sections 4.2 and 5.4 analyze the theoretical time complexity of WinMax and WaterSeeker, respectively.
-This section aims to offer a more comprehensive analysis based on experimental results, taking into account the constant factors preceding the complexity terms.
-Additionally, it will investigate the reasons behind the significant disparity in execution time of WinMax between KGW and Aar.
-In this experimental setting, let the time required to compute a score in the score list be $t_{1}$, and the time required to compute the watermark detection statistic within a local window be $t_{2}$.
-The time expenditure for the WinMax [^2] The WinMax algorithm, in practical implementation, also requires computing the score list for the entire text first, then traversing the windows to calculate the detection statistic.
-algorithm to detect a sample can be expressed as $N\cdot t_{1}+N\cdot(W_{max}-W_{min})\cdot t_{2}$.
-For WaterSeeker, the first step of localization consists of three sub-steps: score list computation, anomaly detection, and fragment connection.
-Each of these sub-steps has a complexity of $O(N)$.
-If we denote the time required for the latter two sub-steps as $t_{3}\cdot N$, then the total time for the first step is $N\cdot t_{1}+N\cdot t_{3}$.
-The second step, local traverse detection, requires $W^{2}\cdot t_{2}$ time.
-Consequently, the total time expenditure for the entire process can be expressed as $N\cdot t_{1}+N\cdot t_{3}+W^{2}\cdot t_{2}$.
-For the Aar (Aaronson & Kirchner, 2022) watermarking method, calculating the watermark detection statistic involves computing a GammaTransform function, which is a relatively complex computation, resulting in a larger value for $t_{2}$.
-In this case, the time difference between WinMax and WaterSeeker will become significant due to the large disparity in the coefficients preceding $t_{2}$.
-In this experiment, with $T=10,000,W_{max}=400,W_{min}=100,W^{2}=2,500$, the coefficient difference amounts to a factor of 1,200.
-For the KGW (Kirchenbauer et al., 2023) watermarking method, calculating the z-score within a local window is a relatively straightforward computation.
-In this case, the time difference between WinMax and WaterSeeker will be reduced.
-Through the above analysis, it is clear that WaterSeeker is a general watermark detection method with consistently low time expenditure across various watermarking algorithms.
-In contrast, WinMax exhibits extremely high time costs in situations where calculating the watermark detection statistic is complex, making it challenging for practical use.
-6.5 Ablation Study
+6.1 Experiment Settings
+Watermarking Methods and Language Models: We selected two representative watermarking algorithms, KGW (Kirchenbauer et al., 2023) and Aar (Aaronson and Kirchner, 2022), each at three strength levels.
+KGW’s strength was set by the $\delta$ parameter (2.0=strong, 1.5=medium, 1.0=weak), while Aar’s strength used the temperature parameter (0.3=strong, 0.2=medium, 0.1=weak).
+We used Llama-2-7b (Touvron et al., 2023) and Mistral-7b (Jiang et al., 2024) as generation models.
+Dataset Construction: The first 30 tokens of each entry in the C4 dataset (Raffel et al., 2020) were used for prompts.
+Watermarked segments of random length (100 to 400 tokens) were then generated using randomly selected watermark strengths.
+For positive examples, one such segment was randomly inserted into each 10,000-token Wikipedia passage (Foundation, ).
+Negative examples consist of unmodified 10,000-token Wikipedia corpus.
+Based on this procedure, four datasets were created, each containing 300 positive and 300 negative examples: KGW-llama, KGW-mistral, Aar-llama, and Aar-mistral.
+The specific distributions of watermark strengths and lengths are detailed in Appendix F.
+Baselines: As introduced in Section 4, Full-text Detection and WinMax (Kirchenbauer et al., 2024) are chosen, as well as Fix-Length Sliding Window method with $W$ of 100, 200, 300 and 400.
+Hyper-parameters: The parameters related to WaterSeeker are as follows: $W=50,k=20$, with a tolerance for fragment connection set to 100.
+The parameter $\theta_{1}$ is set to 0.5, and $\theta_{2}$ is set to 1.5 for both algorithms.
+The threshold selection within the specified window is detailed in Appendix D.
+Notably, careful threshold selection is crucial for maintaining an acceptable false positive rate, as traversing long texts is prone to accumulating false positives.
+Evaluation-related parameters: $\theta$ mentioned in Section 3.2 is set to 0.5.
 Table 3: This table presents the contributions of the first step of WaterSeeker: Suspicious Segment Localization.
@@
 Average Offset (End) 25.985 25.299 28.030 29.459 33.059 30.304
-In this section, we will analyze the effectiveness of the two stages of WaterSeeker through an ablation study.
-The first stage of WaterSeeker, Suspicious Segment Localization, aims to achieve significant coverage of the ground truth segments while keeping the start and end offsets within a specified window size.
-This ensures that subsequent local traversals can reach the gold index.
-Table 3 illustrates the performance of Step 1, showing that the localization results of Suspicious Segment Localization achieve an average coverage level exceeding 0.95 across various watermark algorithms and strengths.
-Additionally, both the average start offset and average end offset are less than 50, remaining within the designated window size.
-This indicates that Step 1 fulfills its intended purpose effectively.
-Local Traverse Detection performs a localized traversal based on the segments narrowed down in the first step, allowing for more refined verification within the window.
-Table 4 illustrates the differences in detection outcomes when using Local Traverse compared to directly applying full-text detection with the localization results.
-It can be found that across different LLMs and watermarking algorithms, Local Traverse consistently enhances the detection performance, making it an indispensable component of WaterSeeker.
-Table 4: This table compares the detection performance of using Local Traverse versus not using it across the four datasets.
+Table 4: Comparison of detection performance with and without Local Traverse Detection across four datasets.
 Model Setting KGW Aar
-FPR FNR F1 IoU FPR FNR F1 IoU
-Llama-2-7b w. local traverse 0.017 0.303 0.813 0.629 0.010 0.433 0.719 0.510
+FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$ FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$
+Llama-2-7b w. local traverse 0.017 0.313 0.806 0.624 0.010 0.440 0.713 0.507
 w/o local traverse 0.003 0.390 0.756 0.520 0.000 0.450 0.710 0.462
-Mistral-7b w. local traverse 0.007 0.287 0.829 0.642 0.010 0.390 0.753 0.542
+Mistral-7b w. local traverse 0.007 0.290 0.827 0.641 0.010 0.390 0.753 0.542
 w/o local traverse 0.003 0.380 0.763 0.534 0.003 0.413 0.738 0.490
-7 Further Application
-Figure 4: An illustration of WaterSeeker applied in AI detection system.
-WaterSeeker not only provides accurate detection results, but its localization capabilities also contribute to building a more transparent and interpretable AI detection system.
-As illustrated in Figure 4, the AI detection system powered by WaterSeeker receives documents for analysis and outputs three components: (1) Detection Result, which indicates whether AI assistance was utilized; (2) Suspicious AI-generated Segments, which highlights segments identified as potentially AI-generated; and (3) AI Ratio, which displays the proportion of content attributed to AI.
-For example, in the context of academic integrity, highlighting suspicious segments provides actionable evidence for assessments and appeals.
-Additionally, reporting the AI ratio allows for more flexible establishment of academic misconduct standards.
+6.2 Detection Performance Analysis
+In Table 1, we present the detection performance of WaterSeeker alongside various baseline algorithms, evaluated on four datasets.
+Given that full-text detection lacks localization capabilities, its FNR is necessarily 1 under the evaluation rule requiring IoU $>\theta$ for successful detection.
+Therefore, we additionally report the FNR and corresponding F1 score for Full-text Detection without the IoU $>\theta$ constraint, highlighted in red.
+Additional performance results for WaterSeeker under various mixing ratios are presented in Appendix G.
+The results demonstrate that WaterSeeker achieves lower FNR and higher F1 score, significantly surpassing other watermark detection algorithms, including Full-text Detection, FLSW-100, FLSW-200, FLSW-300, and FLSW-400, while being comparable to WinMax.
+Based on the principle that WinMax is guaranteed to reach the gold index by evaluating all possible windows, it represents the upper bound of detection performance.
+However, subsequent experiments show that WinMax has extremely high time complexity (Section 6.4).
+Furthermore, the IoU results in the table reveal that WaterSeeker possesses a significantly higher localization capability compared to the baselines, again comparable to WinMax.
+6.3 Adaptability Analysis
+While the main experiment used four datasets with a mix of watermarked segments of varying strengths and lengths, this section compares the detection capabilities of WaterSeeker and the FLSW method across specific watermark strengths and lengths.
+Figure 3 shows that the performance of different detection algorithms follows a consistent trend: watermarked segments with higher strengths and longer lengths are more easily detected.
+FLSW performs well within a length range close to its window size but shows poor adaptability in other ranges due to its fixed-length nature, which leads to dilution effects when detecting longer or shorter watermarked segments.
+In contrast, WaterSeeker exhibits good adaptability across different length ranges by using anomaly extraction techniques and incorporates a top-k score mechanism to adapt to varying watermark strengths.
+6.4 Time Complexity Analysis
+Table 2 shows the average time spent per sample for WaterSeeker and other baseline algorithms during detection.
+WaterSeeker’s time is comparable to full-text detection and FLSW, and significantly lower than WinMax.
+Sections 4.2 and 5.4 analyze the theoretical time complexity of WinMax and WaterSeeker, respectively.
+This section offers a more comprehensive analysis based on experimental results, considering constant factors and investigating the significant disparity in WinMax’s execution time between KGW and Aar.
+Let $t_{1}$ be the time to compute a score in the score list, and $t_{2}$ be the time to compute the watermark detection statistic within a local window.
+WinMax’s time expenditure to detect a sample is $N\cdot t_{1}+N\cdot(W_{\max}-W_{\min})\cdot t_{2}$.
+For WaterSeeker, the localization step consists of score list computation, anomaly detection, and fragment connection, each with complexity $O(N)$.
+Denoting the time for the latter two sub-steps as $t_{3}\cdot N$, the total time for the first step is $N\cdot t_{1}+N\cdot t_{3}$.
+The local traverse detection step requires $W^{2}\cdot t_{2}$ time.
+Thus, WaterSeeker’s total time expenditure is $N\cdot t_{1}+N\cdot t_{3}+W^{2}\cdot t_{2}$.
+For the Aar (Aaronson and Kirchner, 2022) watermarking method, calculating the watermark detection statistic involves a complex GammaTransform function, resulting in a larger $t_{2}$.
+This leads to a significant time difference between WinMax and WaterSeeker due to the large disparity in coefficients preceding $t_{2}$.
+In this experiment, with $N=10,000,W_{\max}=400,W_{\min}=100,W^{2}=2,500$, the coefficient difference is a factor of 1,200.
+For the KGW (Kirchenbauer et al., 2023) watermarking method, calculating the z-score within a local window is relatively straightforward, reducing the time difference between WinMax and WaterSeeker.
+WaterSeeker is a general watermark detection method with consistently low time expenditure across various watermarking algorithms.
+In contrast, WinMax exhibits extremely high time costs when calculating the watermark detection statistic is complex, making it challenging for practical use.
+6.5 Ablation Study
+We analyze the effectiveness of the two stages of WaterSeeker through an ablation study.
+The first stage, Suspicious Segment Localization, aims to achieve high coverage of the ground truth segments while keeping the start and end offsets within a specified window size.
+This ensures subsequent local traversals can reach the gold index.
+Table 3 shows that Step 1 achieves an average coverage exceeding 0.95 across various watermark algorithms and strengths, with average start and end offsets less than 50, remaining within the designated window size.
+Step 1 fulfills its purpose effectively.
+Local Traverse Detection performs a localized traversal based on the segments from Step 1, allowing for more refined verification within the window.
+Table 4 shows that across different LLMs and watermarking algorithms, Local Traverse consistently enhances detection F1 score and average IoU compared to directly applying full-text detection with the localization results, making it an indispensable component of WaterSeeker.
 7 Conclusion
-In this work, we defined a new scenario for detecting watermarked segments in large documents and established its evaluation metrics.
-We identified the limitations of traditional full-text detection methods in this context and proposed a “first locate, then detect” watermark detection algorithm that utilizes a coarse-to-fine strategy.
+This work introduces a new scenario for detecting watermarked segments in large documents and establishes corresponding evaluation metrics.
+We identified the limitations of full-text detection methods in this context and proposed a “first locate, then detect" watermark detection algorithm that utilizes a coarse-to-fine strategy.
 We validated the detection performance and time complexity of our algorithm through a series of analyses and experiments, demonstrating its ability to effectively balance both aspects.
-For future improvements, we recommend exploring more advanced locating methods based on this concept, which may yield better detection results.
-Appendix A Pseudocode of Detection Baselines
+Future research could explore more advanced locating methods based on this concept to potentially yield improved detection results.
+Limitations
+While our method has demonstrated effectiveness in detecting watermarked segments within large documents, there are still some limitations that need to be addressed in future work.
+First, from an evaluation perspective, due to resource constraints, we only conducted experiments on Llama-2-7B and Mistral-7B models.
+The effectiveness of our method on larger and more powerful models remains to be further verified.
+Second, WaterSeeker’s performance may decrease with very short or weak watermarks.
+Enhancing the sensitivity of WaterSeeker to detect shorter and weaker watermarks is an area for future improvement, which may involve refining the anomaly extraction algorithms or incorporating additional contextual analysis.
+Lastly, parameter selection, such as threshold settings, is crucial and can be challenging in different environments.
+Currently, parameters are manually tuned based on observations, so developing adaptive tuning methods, potentially using machine learning, would enhance WaterSeeker’s robustness and practicality across diverse scenarios.
+Appendix A Details of Representative Watermarking Algorithms
+A.1 KGW
+Watermarking.
+In watermarked text generation, the process for the $t$-th token begins by hashing preceding tokens with a secret key, creating a red-green vocabulary partition where green tokens comprise a fraction $\gamma$.
+Green token logits are then incrementally increased by $\delta$, which can be expressed as follows:
+$l^{\prime}_{t}(y)=\begin{cases}l_{t}(y),&y\in R_{t}\\ l_{t}(y)+\delta,&y\in G_{t}\end{cases}$ (1)
+This subtle modification results in watermarked text exhibiting a higher frequency of green tokens compared to non-watermarked text.
+Detection.
+Detecting a KGW watermark entails computing red-green partitions for each position using preceding tokens and the hash function, then calculating the green token proportion using the z-score:
+$z=\frac{|s|_{G}-\gamma N}{\sqrt{\gamma(1-\gamma)N}}$ (2)
+, where $|s|_{G}$ represents the total count of green tokens in the whole text of length $N$.
+A.2 Aar
+Watermarking.
+When generating the $t$-th token, it first involves hashing the preceding tokens using a secret key to obtain a pseudo vector $u_{t}\sim\text{Uniform}([0,1])^{|V|}$.
+The $t$-th token is determined by
+$\mathop{\arg\max}\limits_{y}\,u_{t}(y)^{1/p_{t}(y)},$ (3)
+where $p$ is the probability vector produced by LLM at the $t$-th step. Let’s perform equivalent transformations on it:
+$\bf{y}$ $=\mathop{\arg\max}\limits_{y}\,u_{t}(y)^{1/p_{t}(y)}$ (4)
+$=\mathop{\arg\max}\limits_{y}\,\frac{1}{p_{t}(y)}\log u_{t}(y)$ (5)
+$=\mathop{\arg\min}\limits_{y}\,\frac{1}{p_{t}(y)}\log\frac{1}{u_{t}(y)}$ (6)
+$=\mathop{\arg\min}\limits_{y}\,\log\frac{1}{p_{t}(y)}+\log\log\frac{1}{u_{t}(y)}$ (7)
+$=\mathop{\arg\max}\limits_{y}\,\log p_{t}(y)-\log\log\frac{1}{u_{t}(y)}$ (8)
+Given that the probabilistic output $p_{t}$ of an LLM is derived from the logits $l_{t}$ through a softmax transformation, and when we additionally consider the sampling temperature $T$, Equation 8 becomes equivalent to:
+$\mathop{\arg\max}\limits_{y}\,\cfrac{l_{t}(y)}{T}+G_{t}(y),$ (9)
+where $l_{t}$ is the logits produced by the LLM, and $G_{t}$ is the Gumbel noise: $G_{t}(y)\sim Gumbel(0,1)$.
+The $Gumbel(0,1)$ distribution is defined as follows: if $u\sim\text{Uniform}(0,1)$, then $-\log(-\log(u))$ follows a $Gumbel(0,1)$ distribution.
+It is evident that the temperature $T$ can be utilized to exert control over the watermark strength.
+As the value of $T$ increases, the influence of Gumbel noise on the sampling process becomes more pronounced, consequently resulting in a stronger watermark.
+Detection.
+Detecting an Aar watermark involves calculating the correlation value between the pseudo vector $u_{t}$ and the corresponding token $y_{t}$ in the text to be examined.
+The correlation value can be expressed as:
+$\log\cfrac{1}{1-u_{t}(y_{t})}.$ (10)
+For the entire text, the statistic value can be expressed as:
+[…]
+where $\Gamma$ is the Gamma Transformation function that converts the sum of correlation values to a p-value.
+Appendix B Pseudocode of WaterSeeker and Detection Baselines
+Pseudocode of WaterSeeker, WinMax and FLSW could be found in Algorithm 3, 1 and 2, respectively.
 Algorithm 1 WinMax Algorithm
@@
 13:end procedure
-Appendix B Detailed Proof for Aar
-B.1 Expected p-value of watermarked text
+Algorithm 3 WaterSeeker Algorithm
+1:procedure SuspiciousRegionLocalization(tokens, $W$, $k$, $\theta_{1}$, $\theta_{2}$)
+2:     Compute score list using sliding window
+3:     Calculate $s_{\text{mean}}$, $s_{\text{top-k-mean}}$, and $s_{\text{std}}$
+4:     Detect anomalies using threshold:
+5:     $s_{\text{mean}}+\max((s_{\text{top-k-mean}}-s_{\text{mean}})\cdot\theta_{1},s_{\text{std}}\cdot\theta_{2})$
+6:     Connect nearby outliers and filter short segments
+7:     return filteredSegments
+8:end procedure
+9:procedure LocalTraverseDetection(tokens, suspiciousRegions, $W$, threshold)
+10:     hasWatermark $\leftarrow$ False
+11:     indices $\leftarrow$ [ ]
+12:     for $(s^{\prime},e^{\prime})$ in suspiciousRegions do
+13:         maxScore = $\infty$- ∞, maxIndice = ()
+14:         for $s\in[s^{\prime},s^{\prime}+W)$ and $e\in(e^{\prime}-W,e^{\prime}]$ do
+15:              if DetectWatermark(tokens[$s:e$])$>$ maxScore then
+16:                  maxScore = DetectWatermark(tokens[$s:e$])
+17:                  maxIndice = $(s,e)$
+18:              end if
+19:         end for
+20:         if maxScore $>$ threshold then
+21:              hasWatermark $\leftarrow$ True
+22:              indices.append(maxIndice)
+23:         end if
+24:     end for
+25:     return hasWatermark, indices
+26:end procedure
+27:procedure WaterSeeker(tokens, $W$, $k$, $\theta_{1}$, $\theta_{2}$, threshold)
+28:     suspiciousRegions $\leftarrow$ SuspiciousRegionLocalization(tokens, $W$, $k$, $\theta_{1}$, $\theta_{2}$)
+29:     hasWatermark, indices $\leftarrow$ LocalTraverseDetection(tokens, suspiciousRegions, W, threshold)
+30:     return hasWatermark, indices
+31:end procedure
+Appendix C Detailed Proof for Aar
+C.1 Expected p-value of watermarked text
 In this sub-section, we aim to analyze how the expected p-value for watermarked text varies with changes in the detection window size $W$.
@@
 The expectation of p-value is:
-$E[\text{p-value}]=E[\text{GammaTransform}(S,W,\text{loc}=0,\text{scale}=1)]$ (9)
+$E[\text{p-value}]=E[\Gamma(S,W,\text{loc}=0,\text{scale}=1)]$ (13)
 Since GammaTransform is non-linear, we cannot directly substitute $E[S]$. However, we can use Jensen’s inequality to obtain an approximation:
-$E[\text{GammaTransform}(S,W,0,1)]\geq\text{GammaTransform}(E[S],W,0,1)$ (10)
-Therefore, we can analyze $\text{GammaTransform}(E[S],W,0,1)$ to obtain a lower bound for the expectation of p-value. Define function $f(W)=\text{GammaTransform}(E[S],W,0,1)$:
-  • When $W\leq L$: $f(W)=\text{GammaTransform}(W\mu_{1},W,0,1)$
-  • When $W>L$: $f(W)=\text{GammaTransform}(L\mu_{1}+(W-L)\mu_{0},W,0,1)$
+$E[\Gamma(S,W,0,1)]\geq\Gamma(E[S],W,0,1)$ (14)
+Therefore, we can analyze $\Gamma(E[S],W,0,1)$ to obtain a lower bound for the expectation of p-value. Define function $f(W)=\Gamma(E[S],W,0,1)$:
+  • When $W\leq L$: $f(W)=\Gamma(W\mu_{1},W,0,1)$
+  • When $W>L$: $f(W)=\Gamma(L\mu_{1}+(W-L)\mu_{0},W,0,1)$
 Analyzing the behavior of $f(W)$:
@@
 While this analysis does not strictly prove that the expectation of p-value is minimized exactly at $W=L$, it strongly suggests that the expectation of p-value is likely to reach its minimum value at or very near $W=L$ [^1] For a more rigorous proof, a deeper mathematical analysis of the GammaTransform function or numerical simulations would be necessary.
-B.2 Calculation of p-threshold to control false positive rate
+C.2 Calculation of p-threshold to control false positive rate
 Similar to the analysis for KGW method, we also need to analyze the constraints on the p-value threshold $p^{*}$ when the false positive rate within the specified window is set to be lower than a target value $\alpha$.
@@
 Solving this equation for $p^{*}$, we get $p^{*}=\alpha$, which is also a constant value for different $W$.
-Appendix C Detail of Threshold Selection Within the Specified Window
+Appendix D Detail of Threshold Selection Within the Specified Window
 A key role of threshold selection is to control the false positive rate.
@@
 Therefore, managing the false positive rate within the detection window is crucial in this scenario.
-In the experiment, we set the targeted false positive rate $\alpha$ within the detection window to $10^{-6}$.
-C.1 Rationale for Setting $\alpha$ to $10^{-6}$
+In the experiment, we set the target false positive rate $\alpha$ within the detection window to $10^{-6}$.
+D.1 Rationale for setting $\alpha$ to $10^{-6}$
 Table 5: Simulated FPR of WaterSeeker using 10,000 samples for each watermarking method. The targeted false positive rate within the detection window is set to $10^{-6}$.
@@
 WaterSeeker, WinMax, and FLSW all involve employing sliding windows for text traversal and conduct full-text detection within each window.
-As these windows overlap, they cannot be treated as independent, making it challenging to derive a theoretical upper bound for the document-level FPR from the targeted FPR within each window.
-Given this, we utilize large-scale data simulation to demonstrate that, with a targeted false positive rate of $10^{-6}$ within each window, our proposed method WaterSeeker maintains an acceptable false positive rate.
+As these windows overlap, they cannot be treated as independent, making it challenging to derive a theoretical upper bound for the document-level FPR from the target FPR within each window.
+Given this, we utilize large-scale data simulation to demonstrate that, with a target false positive rate of $10^{-6}$ within each window, our proposed method WaterSeeker maintains an acceptable false positive rate.
 For the KGW method, we set $\gamma=0.5$ in our experiments, meaning each token in non-watermarked text has a 0.5 probability of being green and 0.5 probability of being red.
@@
 In the simulation, we again generate 10,000 samples, each containing 10,000 tokens, with each token randomly assigned a floating-point number from $[0,1]$.
-We then apply WaterSeeker to detect watermarked segments within these samples, setting the targeted false positive rate within the detection window to $10^{-6}$.
+We then apply WaterSeeker to detect watermarked segments within these samples, setting the target false positive rate within the detection window to $10^{-6}$.
 The large-scale simulation results in Table 5 demonstrate that WaterSeeker maintains a false positive rate of approximately 0.005, which is considered acceptable.
-For scenarios requiring more stringent FPR control, the targeted false positive rate can be adjusted downward.
+For scenarios requiring more stringent FPR control, the target false positive rate can be adjusted downward.
 However, this inevitably compromises the detection rate, highlighting a key challenge in watermarked segment detection within large documents.
-C.2 Setting the Threshold to Achieve a Target False Positive Rate $\alpha$
+D.2 Setting the threshold to achieve a target false positive rate $\alpha$
 For KGW, as analyzed in Section 5.1, when the window size is large, we can approximate using the Central Limit Theorem, resulting in $z^{*}=\Phi^{-1}(1-\alpha)$.
@@
 For segments shorter than 200, we use the binomial distribution and retrieve the corresponding threshold from the pre-computed dictionary.
-For Aar, as analyzed in Appendix B, $p^{*}=\alpha$, therefore set $p^{*}=10^{-6}$ for all lengths.
-Appendix D Sample Distribution in the Main Experiment
+For Aar, as analyzed in Appendix C, $p^{*}=\alpha$, therefore set $p^{*}=10^{-6}$ for all lengths.
+Appendix E Impact of Window Size on Watermark Intensity Calculation
+(a) $W=1$
+(b) $W=10$
+(c) $W=30$
+(d) $W=50$
+Figure 4: Case study: Impact of varying window sizes on watermark intensity calculation in the KGW algorithm.
+(a) $W=1$
+(b) $W=10$
+(c) $W=30$
+(d) $W=50$
+Figure 5: Case study: Impact of varying window sizes on watermark intensity calculation in the Aar algorithm.
+The first step in WaterSeeker is score list computation.
+In this step, selecting an appropriate window size $W$ for calculating mean scores is crucial.
+A small $W$ introduces excessive noise, while a large $W$ reduces granularity and increases computational time due to the need to examine $W^{2}$ windows during local traversal.
+Therefore, we aim to determine an appropriate window size that is relatively small while still providing a sufficiently smooth representation of watermark intensity throughout the text.
+We present a case study comparing watermark intensity calculations using window sizes $W=1,10,30,$ and $50$.
+The analysis encompasses the ground truth segment and 500 tokens on either side.
+Figures 4 and 5 illustrate the results for the KGW and Aar algorithms, respectively.
+The intensity curves reveal that small window sizes, particularly $W\leq 10$, introduce significant fluctuations.
+While $W=30$ exhibits reduced noise, it still presents instabilities, as shown in Figure 5(c) (the ground truth segment part).
+Overall, $W=50$ demonstrates the least noise.
+Consequently, we adopt $W=50$ for our main experiments.
+Appendix F Sample Distribution in the Main Experiment
 It can be observed from Figure 6 that all four datasets include samples of watermarked segments with varying intensities and lengths, demonstrating the comprehensiveness and fairness of the dataset construction.
@@
 Figure 6: Distribution of sample numbers for the four datasets involved in the main experiment.
+Appendix G Further Experimental Results under Different Mixing Ratios
+In main experiment, 100-400 watermarked tokens are inserted into non-watermarked text of 10,000 tokens, resulting in watermarked text mixing ratios of 1% to 4%.
+This section expands the evaluation by testing the performance of various methods across a broader range of mixing ratios, achieved by adjusting the length of the non-watermarked text.
+We deliberately avoid adjusting the mixing ratio by inserting multiple watermark segments into the non-watermarked text, as the WinMax method can only return the highest-scoring segment and is not designed for multi-segment insertion scenarios.
+Additionally, we do not use fixed-length watermark segments because in real-world detection scenarios, the length of the watermark segment is typically unknown.
+Moreover, if the watermark length were fixed and known, a simple Fixed-length Sliding Window detection method (with the window size equal to the watermark segment length) would suffice to solve the problem.
+We provided further experimental results under three different settings:
+  1. 1.
+    100-400 watermarked tokens are inserted into 5,000 non-watermarked tokens, with a mixing ratio of 2% (100 / 5100) -7.4% (400 / 5400). The results are demonstrated in Table 6.
+  2. 2.
+    100-400 watermarked tokens are inserted into 2,000 non-watermarked tokens, with a mixing ratio of 4.8% (100 / 2100) -16.7% (400 / 2400). The results are demonstrated in Table 7.
+  3. 3.
+    100-400 watermarked tokens are inserted into 500 non-watermarked tokens, with a mixing ratio of 16.7% (100 / 600) -44.4% (400 / 900). The results are demonstrated in Table 8.
+Table 6: Comparison of WaterSeeker with other baselines under the mixing ratio of 2%-7.4%.
+Model Method KGW Aar
+FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$ FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$
+Llama-2-7b Full-text Detection 0.000 1.000 [0.990] 0.000 [0.020] 0.001 0.000 1.000 [0.957] 0.000 [0.083] 0.002
+WinMax 0.003 0.317 0.810 0.621 0.010 0.380 0.761 0.565
+FLSW-100 0.000 0.520 0.649 0.413 0.003 0.523 0.644 0.375
+FLSW-200 0.000 0.457 0.704 0.404 0.000 0.560 0.611 0.350
+FLSW-300 0.003 0.663 0.502 0.297 0.000 0.767 0.378 0.283
+FLSW-400 0.003 0.867 0.235 0.230 0.000 0.920 0.148 0.205
+WaterSeeker(Ours) 0.003 0.357 0.781 0.570 0.010 0.407 0.740 0.536
+Mistral-7b Full-text Detection 0.000 1.000 [0.990] 0.000 [0.020] 0.000 0.000 1.000 [0.937] 0.000 [0.119] 0.004
+WinMax 0.003 0.227 0.871 0.700 0.010 0.370 0.768 0.570
+FLSW-100 0.000 0.503 0.664 0.443 0.000 0.543 0.627 0.364
+FLSW-200 0.000 0.350 0.788 0.497 0.003 0.540 0.629 0.362
+FLSW-300 0.000 0.527 0.643 0.392 0.003 0.763 0.382 0.269
+FLSW-400 0.000 0.783 0.357 0.281 0.003 0.910 0.165 0.203
+WaterSeeker(Ours) 0.000 0.243 0.861 0.676 0.010 0.383 0.758 0.548
+Table 7: Comparison of WaterSeeker with other baselines under the mixing ratio of 4.8%-16.7%.
+Model Method KGW Aar
+FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$ FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$
+Llama-2-7b Full-text Detection 0.000 1.000 [0.967] 0.000 [0.065] 0.005 0.000 1.000 [0.873] 0.000 [0.225] 0.018
+WinMax 0.003 0.257 0.851 0.669 0.003 0.413 0.738 0.533
+FLSW-100 0.000 0.473 0.690 0.440 0.003 0.570 0.600 0.350
+FLSW-200 0.000 0.420 0.734 0.446 0.000 0.607 0.565 0.327
+FLSW-300 0.003 0.630 0.539 0.339 0.000 0.777 0.365 0.257
+FLSW-400 0.003 0.850 0.260 0.246 0.000 0.883 0.209 0.200
+WaterSeeker(Ours) 0.003 0.287 0.831 0.642 0.003 0.443 0.714 0.501
+Mistral-7b Full-text Detection 0.000 1.000 [0.980] 0.000 [0.039] 0.003 0.000 1.000 [0.870] 0.000 [0.230] 0.017
+WinMax 0.000 0.220 0.876 0.710 0.003 0.353 0.783 0.590
+FLSW-100 0.000 0.470 0.693 0.447 0.000 0.487 0.678 0.415
+FLSW-200 0.000 0.360 0.780 0.497 0.003 0.500 0.665 0.371
+FLSW-300 0.000 0.533 0.636 0.386 0.003 0.760 0.386 0.285
+FLSW-400 0.000 0.747 0.404 0.288 0.003 0.887 0.203 0.212
+WaterSeeker(Ours) 0.000 0.227 0.872 0.701 0.003 0.363 0.776 0.574
+Table 8: Comparison of WaterSeeker with other baselines under the mixing ratio of 16.7%-44.4%.
+Model Method KGW Aar
+FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$ FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$
+Llama-2-7b Full-text Detection 0.000 1.000 [0.787] 0.000 [0.352] 0.081 0.000 1.000 [0.693] 0.000 [0.469] 0.109
+WinMax 0.000 0.247 0.859 0.685 0.000 0.373 0.770 0.568
+FLSW-100 0.000 0.470 0.693 0.445 0.000 0.557 0.614 0.380
+FLSW-200 0.000 0.383 0.763 0.461 0.000 0.553 0.618 0.340
+FLSW-300 0.000 0.553 0.618 0.360 0.000 0.690 0.473 0.298
+FLSW-400 0.000 0.740 0.413 0.265 0.000 0.837 0.281 0.231
+WaterSeeker(Ours) 0.000 0.277 0.839 0.647 0.000 0.410 0.742 0.522
+Mistral-7b Full-text Detection 0.000 1.000 [0.847] 0.000 [0.266] 0.061 0.000 1.000 [0.697] 0.000 [0.465] 0.112
+WinMax 0.000 0.270 0.844 0.652 0.000 0.350 0.788 0.588
+FLSW-100 0.000 0.510 0.658 0.432 0.000 0.530 0.639 0.393
+FLSW-200 0.000 0.407 0.745 0.442 0.000 0.510 0.658 0.369
+FLSW-300 0.000 0.570 0.601 0.352 0.000 0.690 0.473 0.294
+FLSW-400 0.000 0.760 0.387 0.243 0.000 0.827 0.295 0.246
+WaterSeeker(Ours) 0.000 0.303 0.821 0.612 0.000 0.407 0.745 0.543
+Experimental results demonstrate that WaterSeeker exhibits high robustness across various mixing ratios, showing consistent performance trends across different settings.
+Specifically, WaterSeeker significantly outperforms full-text detection and FLSW methods in both detection accuracy and localization effectiveness, while achieving comparable performance to the upper bound method, WinMax.
+Appendix H Further Application Scenario
+Figure 7: An illustration of WaterSeeker applied in AI detection system.
+WaterSeeker not only provides accurate detection results, but its localization capabilities also contribute to building a more transparent and interpretable AI detection system.
+As illustrated in Figure 7, the AI detection system powered by WaterSeeker receives documents for analysis and outputs three components: (1) Detection Result, which indicates whether AI assistance was utilized; (2) Suspicious AI-generated Segments, which highlights segments identified as potentially AI-generated; and (3) AI Ratio, which displays the proportion of content attributed to AI.
+For example, in the context of academic integrity, highlighting suspicious segments provides actionable evidence for assessments and appeals.
+Additionally, reporting the AI ratio allows for more flexible establishment of academic misconduct standards.

### v4

@@
 Theoretical analysis and experimental results demonstrate that WaterSeeker achieves a superior balance between detection accuracy and computational efficiency.
-Moreover, WaterSeeker’s localization ability supports the development of interpretable AI detection systems.
-This work pioneers a new direction in watermarked segment detection, facilitating more reliable AI-generated content identification.
+Moreover, its localization capability lays the foundation for building interpretable AI detection systems.
 Our code is available at https://github.com/THU-BPM/WaterSeeker.
 WaterSeeker: Pioneering Efficient Detection of Watermarked Segments in Large Documents
-Leyi Pan1,     Aiwei Liu 1,     Yijian Lu1,     Zitian Gao2,     Yichen Di1,     Lijie Wen1thanks: Corresponding author,     Irwin King3,     Philip S.
-Yu4 1Tsinghua University    2The University of Sydney 3The Chinese University of Hong Kong    4University of Illinois at Chicago panly24@mails.tsinghua.edu.cn, liuaw20@mails.tsinghua.edu.cn, wenlj@tsinghua.edu.cn
+Leyi Pan1thanks: Work was done during the intern at Zhipu AI.,     Aiwei Liu 1,     Yijian Lu1,     Zitian Gao2,     Yichen Di1,     Shiyu Huang3,     Lijie Wen1thanks: Corresponding author,     Irwin King4,     Philip S.
+Yu5 1Tsinghua University    2The University of Sydney    3Zhipu AI 4The Chinese University of Hong Kong    5University of Illinois at Chicago panly24@mails.tsinghua.edu.cn, liuaw20@mails.tsinghua.edu.cn, wenlj@tsinghua.edu.cn
 1 Introduction
-As large language models (LLMs) generate high-quality text, they address practical challenges but also raise concerns such as misinformation (Liu et al., 2023b; Chen and Shu, 2024) and copyright infringement (Rillig et al., 2023).
+As large language models (LLMs) generate high-quality text, they address practical challenges but also raise concerns such as misinformation (Megías et al., 2022; Chen and Shu, 2024) and copyright infringement (Rillig et al., 2023).
 LLM watermarking technology has emerged to tackle these issues by embedding specific information (watermarks) during text generation, allowing for accurate detection through specialized algorithms.
-Existing watermarking methods primarily focus on distinguishing fully watermarked text from non-watermarked content (Kirchenbauer et al., 2023; Zhao et al., 2024; Liu et al., 2024b; Lu et al., 2024; Aaronson and Kirchner, 2022; Kuditipudi et al., 2024).
-However, in real-world scenarios, LLMs often generate only brief segments within longer documents, making it essential to detect these watermarked sections among large amounts of non-watermarked text.
-Most previous algorithms relied on computing statistics across the entire document for detection, which we refer to as full-text detection methods (detailed in Section 2).
-However, these full-text detection methods struggle when small watermarked sections are mixed with large volumes of natural text due to dilution effects, as illustrated in Figure 1(a).
-To the best of our knowledge, the only existing work addressing this scenario is the WinMax (Kirchenbauer et al., 2024) algorithm, which examines all possible window sizes and traverses the full text for each size.
-However, this method suffers from high time complexity.
-Additionally, our self-constructed baseline, the Fixed-Length Sliding Window (FLSW) method, faces the issue of watermark strength dilution due to inflexible window sizes, which lowers detection effectiveness.
-(a) Full-text detection methods often struggle in watermarked segment detection scenarios due to the dilution effect.
-(b) An illustration of the WaterSeeker algorithm, using “first locate, then detect" strategy to balance time complexity and detection performance.
-Figure 1: An overview of watermark detection scenarios and the proposed WaterSeeker algorithm.
-To address these issues, we propose a novel watermark detection method called WaterSeeker.
-WaterSeeker follows a "first locate, then detect" approach, as shown in Figure 1(b).
+Current watermark detection methods (Kirchenbauer et al., 2023; Zhao et al., 2024; Liu et al., 2024; Aaronson and Kirchner, 2022; Lu et al., 2024; Lee et al., 2023; Hu et al., 2024; Wu et al., 2023) first calculate watermark scores for individual tokens through single token detection, then compute statistics across the entire document for classification.
+While these full-text detection methods effectively distinguish between fully watermarked and non-watermarked texts, they fail in real-world scenarios where LLMs generate only brief segments within longer documents, due to dilution effects, as shown in Figure 1.
+To the best of our knowledge, the WinMax algorithm (Kirchenbauer et al., 2024) is the only work addressing this limitation by examining all possible window sizes and selecting the maximum statistical score across all windows as the final detection result, but suffers from high time complexity.
+Figure 1: While full-text detection methods effectively differentiate between fully watermarked and non-watermarked texts, they often struggle with watermarked segment detection due to the dilution effect.
+To address this, WaterSeeker employs a "first locate, then detect" strategy, which narrows the detection range before conducting local traversal for further verification.
+To address these issues, we propose a novel and general watermark detection method called WaterSeeker.
+WaterSeeker follows a "first locate, then detect" approach, as shown in Figure 1.
 It initially employs a low-complexity anomaly points extraction algorithm to identify suspected watermark regions, narrowing the detection target from a long text to a small segment encompassing the ground truth segment.
@@
 WaterSeeker significantly outperformed the baselines in balancing time complexity and detection performance.
-We also assessed its adaptability to different watermark strengths and segment lengths, and evaluated its two-stage utility through an ablation study.
+Moreover, it demonstrates good adaptability to different watermark strengths, segment lengths, and document lengths, while achieving robust performance against text edit attacks.
 In summary, the contributions of this work are as follows:
@@
   • WaterSeeker outperforms baselines in achieving a superior balance between time complexity and detection effectiveness.
-  • WaterSeeker not only provides accurate detection results but also pinpoints watermark locations precisely, helping to create more explainable AI detection systems (detailed in Appendix H).
+  • Further experiments demonstrate that WaterSeeker exhibits strong adaptability across various watermark strengths, segment lengths, and document lengths, while also being robust against text editing attacks.
 2 Related Work
-Currently, mainstream LLM watermarking methods involve modifying the inference phase by altering logits or influencing token sampling (Liu et al., 2023a; Pan et al., 2024; Liu et al., 2024c).
-The KGW family (Kirchenbauer et al., 2023; Zhao et al., 2024; Hu et al., 2024; Liu et al., 2024b; He et al., 2024; Lu et al., 2024; Liu et al., 2024a) categorizes vocabulary into green and red lists, biasing towards green tokens during generation.
+Currently, mainstream LLM watermarking methods involve modifying the inference phase by altering logits or influencing token sampling (Liu et al., 2023; Pan et al., 2024).
+The KGW family (Kirchenbauer et al., 2023; Zhao et al., 2024; Hu et al., 2024; Liu et al., 2024; Wu et al., 2023) categorizes vocabulary into green and red lists, biasing towards green tokens during generation.
 The bias value is typically determined by the parameter $\delta$, which reflects the watermark strength.
-Detection involves calculating the z-score of green tokens among the entire document; exceeding a threshold indicates watermarking.
+For these methods, single token detection determines whether each token belongs to the green list, while full-text detection involves calculating the z-score of green tokens across the entire document; exceeding a threshold indicates watermarking.
 On the other hand, the Aar family (Aaronson and Kirchner, 2022; Christ et al., 2024; Kuditipudi et al., 2024) uses pseudo-random sequences to guide token sampling.
@@
 Watermark strength is controlled by sampling temperature.
-For detection, it also employs global statistics: it sums the correlation values of each token with the pseudo-random vector and then performs gamma transformation to derive the detection confidence.
+In these methods, single token detection calculates the correlation value between each token and $u$, while full-text detection applies gamma transformation to derive the detection confidence.
 Details of the KGW and Aar watermarking algorithms can be found in Appendix A.
-Despite the high accuracy of watermarking algorithms for detecting LLM-generated text, previous studies have primarily focused on distinguishing between fully watermarked and non-watermarked text, overlooking the possibility that LLMs may only generate small segments within large documents.
-In such cases, watermark detection algorithms based on full-text statistics can fail due to the dilution effect.
+Despite the high accuracy of watermarking algorithms for distinguishing between fully watermarked and non-watermarked text, their performance fall sharply when detecting watermarked segments within large documents.
 A few studies have mentioned copy-paste attack (Kirchenbauer et al., 2024; Yoo et al., 2023; Wang et al., 2024), which involves mixing a portion of watermarked text with non-watermarked content, similar to our scenario.
@@
 However, as they did not develop specific detection mechanisms for this situation, their findings showed that their methods were not robust against this type of attack.
-In existing studies, WinMax (Kirchenbauer et al., 2024) is the only watermark detection algorithm specifically designed for this scenario.
-It examines all possible window sizes and traverses the entire text for each size to identify the maximum local score for threshold comparison.
-While this method shows some effectiveness, its high time complexity renders it impractical for real-world applications.
-Therefore, to achieve effective detection of watermarked segments inserted into large documents while maintaining controllable time complexity, we propose WaterSeeker, a novel and general method that uses “first locate then detect" strategy.
+Among existing studies, only WinMax (Kirchenbauer et al., 2024) specifically addresses watermarked segment detection by examining all possible window sizes to find the highest local statistics.
+However, its high time complexity limits practical application.
+To address this limitation, we propose WaterSeeker, a novel method that employs a "first locate then detect" strategy to achieve efficient detection of watermarked segments in large documents.
 3 Problem Formulation
-3.1 Definition
-The problem of detecting watermarked segments in a long document is defined as follows: Let $N$ be a long text with a randomly inserted watermarked segment of length $L$.
-We denote the starting index of the watermarked segment in $N$ as $s$ and the ending index as $e$, such that: $L=e-s+1$.
-The objective is to determine the presence and location of a watermarked segment in $N$.
-This can be framed as a binary classification problem with additional localization, where the detection algorithm outputs:
+Definition.
+Given a text of length $N$ containing $m$ watermarked segments at position $[(s_{1},e_{1}),...(s_{m},e_{m})]$, the objective is to determine the presence and location of the watermarked segment.
+The detection algorithm outputs: {‘has_watermark’: True/False, ‘indices’: $[(s_{1}^{\prime},e_{1}^{\prime}),...,(s_{m^{\prime}}^{\prime},{e}_{m^{\prime}}^{\prime})]$}.
 […]
-[…]
-Here, ‘has_watermark’ is a boolean value indicating the presence of a watermark, and ‘indices’ is a list of pairs $(s^{\prime},e^{\prime})$ representing the start and end indices of detected watermarked segments.
-3.2 Evaluation
-[…]
-$\text{IoU}=\cfrac{L_{\text{intersection}}}{L_{\text{union}}}>\theta.$
-Based on this criteria, we will report the following metrics of the binary classification result: False Positive Rate (FPR), False Negative Rate (FNR), and F1 Score.
-Moreover, we will include the average IoU between the detected segments and ground truth segments as a supplementary metric, further demonstrating the accuracy of the localization.
+$\text{IoU}=\cfrac{L_{\text{intersection}}}{L_{\text{union}}}>0.$ (1)
+Based on these criteria, we evaluate the classification performance using False Positive Rate (FPR), False Negative Rate (FNR) and F1 Score, as well as the localization performance using average IoU between detected and ground truth segments.
 4 Baseline Methods
-4.1 Full-text Detection
-[…]
-4.2 WinMax
-WinMax (Kirchenbauer et al., 2024) involves iterating through all possible window sizes, and for each window size, the entire text is traversed to calculate the z-score for each local window, taking the maximum z-score and comparing it against a specified threshold.
-In practical applications, a minimum window size $W_{min}$ and a maximum window size $W_{max}$ are typically set to control the number of iterations during the traversal.
+Full-text Detection. As explained in Section 1, involves calculate watermark scores for individual tokens and compute statistics across the entire documents for classification.
+WinMax (Kirchenbauer et al., 2024) involves iterating through all possible window sizes, and for each window size, the entire text is traversed to compute statistics for each local window, taking the maximum score for final results.
 The detection process can be described by the following formula:
-$z_{\text{win-max}}=\max_{w\in[W_{min},W_{max}]}\left(\max_{i}\left(\frac{|s|_{G,i}-\gamma w}{\sqrt{\gamma(1-\gamma)w}}\right)\right),$
-where $w$ is the length of the local window, and $|s|_{G,i}$ represents the count of green tokens within the $i$-th local window of length $w$.
-The time complexity of WinMax is $O((W_{max}-W_{min})\times N)$, where $N$ is the total length of the text.
-Due to the uncertain length of the watermarked segment inserted into the document, if there is a large difference between $W_{max}$ and $W_{min}$, the worst-case complexity can reach $O(N^{2})$.
-4.3 Fix-Length Sliding Window
-Fix-Length Sliding Window (FLSW) is a self-constructed method that uses a fixed-length window to traverse the text, calculating statistics within each local window.
-The text is flagged as watermarked if any statistic score exceeds the threshold.
-For localization, the method records each pair of indices $(s,e)$ meeting the condition and uses a contiguous segment merging method with tolerance.
-While straightforward, FLSW struggles with varying watermark lengths due to its fixed-length nature, leading to dilution effects.
-Section 5.1 analyzes this effect using KGW (Kirchenbauer et al., 2023) and Aar (Aaronson and Kirchner, 2022) as examples, providing theoretical support for WaterSeeker.
+$\text{score}=\max_{1\leq w\leq N}\max_{1\leq i\leq N-w+1}F(x_{i:i+w-1}),$ (2)
+where $w$ is the length of the local window, $w$ represents the text tokens, and $F$ is the statistical function.
+The time complexity is evidently $O(N^{2})$.
+We also introduce a WinMax variant where window size increases by intervals $>1$, reducing complexity to $O(N^{2}/\text{interval})$.
+Fix-Length Sliding Window (FLSW) is a self-constructed method that uses a fixed-length window to traverse the text.
+The text is flagged as watermarked if any statistic score within the local windows exceeds the threshold.
+The pseudocode for all baselines can be found at Appendix C.
 5 Proposed Method: WaterSeeker
 5.1 Theoretical Basis: Gold Index is the Best
+Figure 2: Expected z-score and the corresponding threshold $z^{*}$ across various $W$, $\alpha=10^{-6},\gamma=0.5,\gamma_{1}=0.75$.
 This section provides the theoretical foundation of WaterSeeker, showing that using actual start and end indices (gold index) for watermark detection achieves the highest expected detection rate.
-We use KGW and Aar as examples, since most watermarking algorithms are slight modifications of these two methods.
-Thus, analyzing these approaches offers broad applicability.
-For the KGW method, assuming $\gamma_{1}>\gamma$ is the proportion of green tokens in the watermarked part.
-Let’s analyze the effect of window size $W$ on this statistic: (1) When $W<L$ (window size smaller than watermark length):
-$E[z_{W}]=\frac{W\gamma_{1}-\gamma W}{\sqrt{\gamma(1-\gamma)W}}=\sqrt{W}\cdot\frac{\gamma_{1}-\gamma}{\sqrt{\gamma(1-\gamma)}}.$
-(2) When $W>L$ (window size larger than watermark length):
-$E[z_{W}]=\frac{L\gamma_{1}+(W-L)\gamma-\gamma W}{\sqrt{\gamma(1-\gamma)W}}=\frac{L(\gamma_{1}-\gamma)}{\sqrt{\gamma(1-\gamma)W}}.$
+We analyze using KGW (Kirchenbauer et al., 2023) as a case study.
+Assuming $\gamma_{1}>\gamma$ is the proportion of green tokens in the watermarked part, the watermark segment has a length $L$, and the statistical function $F$ is the z-score compuatation (detailed in Appendix A).
+Let’s analyze the effect of window size $W$ on this statistic: (1) When $W<L$:
+$E[z_{W}]=\frac{W\gamma_{1}-\gamma W}{\sqrt{\gamma(1-\gamma)W}}=\sqrt{W}\cdot\frac{\gamma_{1}-\gamma}{\sqrt{\gamma(1-\gamma)}}.$ (3)
+(2) When $W>L$:
+$E[z_{W}]=\frac{L\gamma_{1}+(W-L)\gamma-\gamma W}{\sqrt{\gamma(1-\gamma)W}}=\frac{L(\gamma_{1}-\gamma)}{\sqrt{\gamma(1-\gamma)W}}.$ (4)
 From this, we can conclude that when $W=L$, the z-score reaches its maximum.
-During detection, we aim for a higher z-score for positive cases while setting an appropriate threshold to balance the false positive rate.
-Next, we will analyze the constraints on the z-threshold $z^{*}$ when the false positive rate within the specified window is set to be lower than a target value $\alpha$.
-We start with the binomial distribution $B(W,\gamma)$ that describes the number of green tokens in a window of size $W$.
-For large sample sizes (typically $W>50$), we can approximate this as $|s|_{G}\sim N(W\gamma,W\gamma(1-\gamma))$.
-Therefore, we have $z\sim N(0,1)$.
-The false positive rate $\alpha$ represents the area in the right tail of this distribution beyond $z^{*}$.
-Mathematically, this is expressed as $\alpha=1-\Phi(z^{*})$.
-Solving for $z^{*}$, we get: $z^{*}=\Phi^{(-1)}(1-\alpha)$, which is a constant value for different $W$.
-For the Aar method, the p-value is calculated by applying a gamma transformation to the sum of correlation values of all the tokens:
-[…]
-We denote $\mu=E[\log(\frac{1}{1-u_{i}})]$ as the expected correction value for a single token.
-Similar to KGW, we assume $\mu_{1}>\mu$ represents the expected correlation value for the watermarked portion.
-Due to the complexity of the Gamma-Transformation, the detailed proof can be found in Appendix C.
-Based on the above proof, we simulated the expected score of the statistic and the corresponding threshold as they vary with $W$ for $L=200$ using real data, as shown in Figure 2.
-It demonstrates that when the window size corresponds to the gold index, the highest detection rate can be achieved while maintaining an acceptable FPR.
-Therefore, the watermark dilution effect refers to the situation where the window used for detection is not optimal, resulting in the computed statistics not reaching their peak or minimum, which in turn affects detection performance.
-(a) KGW, $\alpha=10^{-6},\gamma=0.5,\gamma_{1}=0.75$
-(b) Aar, $\alpha=10^{-6},\mu=1.0,\mu_{1}=1.6$
-Figure 2: Expected score of the statistics and the corresponding threshold across various $W$.
+During detection, we aim to maximize positive case z-scores while maintaining false positive rate below a target threshold $\alpha$.
+For a window of size $W$, the number of green tokens follows $B(W,\gamma)$, which approximates to $N(W\gamma,W\gamma(1-\gamma))$ for large $W$.
+This yields $z\sim N(0,1)$, and the threshold $z^{*}$ is derived as: $z^{*}=\Phi^{(-1)}(1-\alpha)$, which is a constant value for different $W$.
+The simulation result using real data are shown in Figure 2.
 5.2 Suspicious Region Localization
-Based on the theoretical analysis presented in Section 5.1, the design concept of WaterSeeker employs a coarse-to-fine process to gradually approximate the gold index, thereby achieving high detection performance.
-In the first step (coarse step), WaterSeeker uses a localization algorithm to identify suspicious watermarked regions.
-This process narrows the detection target to a small segment that encompasses the ground truth segment, while ensuring relatively small deviation from the gold index.
-Below is the detailed process of the localization algorithm:
-(1) Score List Computation: A small sliding window (i.e., $W=50$) is used to traverse the entire text.
-For each window position, a statistical measure is computed to reflect the average watermark intensity.
-In the KGW method, this statistic is $|s|_{G}/W$, representing the density of green tokens within the window.
-In the Aar method, it is calculated as $\frac{\sum\log(1/(1-u))}{W}$, representing the average correlation value between tokens and the pseudo-random sequence within the window.
-This results in a score list $s$ of length $N-W+1$, where $s_{i}$ represents the average watermark intensity from $n_{i}$ to $n_{i}+W$.
-(2) Anomaly Extraction: In this step, we design an anomaly extraction algorithm to identify outliers from the score list.
-We calculate the mean of the score list, the top-k mean, and the standard deviation, denoted as $s_{\text{mean}}$, $s_{\text{top-k-mean}}$, and $s_{\text{std}}$, respectively.
-The extracted outliers are the points that satisfy the following condition, where $\theta_{1}$ and $\theta_{2}$ are hyperparameters:
-$\text{score}>s_{\text{mean}}+\max((s_{\text{top-k-mean}}-s_{\text{mean}})\cdot\theta_{1},s_{\text{std}}\cdot\theta_{2}).$
-In samples containing a watermarked segment, the first term in the max expression becomes significant.
-Introducing the top-k mean allows for adaptive adjustment of the deviation from the mean based on different watermark strengths.
-Conversely, in non-watermarked samples, the difference between the top-k mean and the mean is relatively small.
-In this case, to reduce false positives, a point must deviate from the mean by at least $s_{\text{std}}\cdot\theta_{2}$ to be considered an anomaly.
-(3) Fragment Connection: In this step, we use a connecting method with a certain tolerance to link nearby outliers, then filter out segments that are too short, returning a list of indices.
-Table 1: We compared the detection performance of WaterSeeker with other methods, including Full-text Detection, WinMax (Kirchenbauer et al., 2024), and FLSW.
-The reported metrics include FPR, FNR, F1 score, and the average IoU between detected segments and ground truth segments in positive samples.
-Model Method KGW Aar
-FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$ FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$
-Llama-2-7b Full-text Detection 0.000 1.000 [0.987] 0.000 [0.026] 0.000 0.000 1.000 [0.967] 0.000 [0.065] 0.001
-WinMax 0.017 0.273 0.834 0.661 0.017 0.410 0.734 0.537
-FLSW-100 0.003 0.473 0.688 0.448 0.003 0.547 0.622 0.363
-FLSW-200 0.003 0.440 0.716 0.417 0.000 0.577 0.595 0.338
-FLSW-300 0.007 0.683 0.479 0.313 0.000 0.783 0.356 0.265
-FLSW-400 0.003 0.897 0.187 0.230 0.003 0.933 0.125 0.191
-WaterSeeker(Ours) 0.017 0.313 0.806 0.624 0.010 0.440 0.713 0.507
-Mistral-7b Full-text Detection 0.000 1.000 [0.990] 0.000 [0.020] 0.000 0.000 1.000 [0.973] 0.000 [0.052] 0.001
-WinMax 0.010 0.277 0.835 0.656 0.013 0.370 0.767 0.562
-FLSW-100 0.000 0.500 0.667 0.433 0.000 0.497 0.670 0.411
-FLSW-200 0.000 0.410 0.742 0.452 0.003 0.560 0.610 0.359
-FLSW-300 0.003 0.577 0.593 0.341 0.003 0.827 0.295 0.254
-FLSW-400 0.003 0.810 0.318 0.244 0.007 0.943 0.107 0.195
-WaterSeeker(Ours) 0.007 0.290 0.827 0.641 0.010 0.390 0.753 0.542
-The scores within watermarked segments are relatively stable and close to the top-k mean, while scores outside the watermark remain stable and close to the overall mean due to the dilution effect of large natural text.
-When using $(s_{\text{top-k-mean}}-s_{\text{mean}})\cdot\theta_{1}\quad(\theta_{1}<1)$ as a distinction criterion, the extracted abnormal segment’s start and end points ($s^{\prime}$ and $e^{\prime}$) generally satisfy $s^{\prime}\in(s-W,s)$ and $e^{\prime}\in(e,e+W)$.
-This ensures that the extracted suspicious watermarked regions likely cover the actual segments, with starting and ending deviations within a window size.
+Based on the theoretical analysis, WaterSeeker employs a coarse-to-fine process to gradually approximate the gold index.
+In the coarse step, a localization algorithm identifies suspicious regions and narrows detection to a small segment containing the ground truth, while maintaining minimal deviation.
+This step involves three sub-steps:
+(1) Score List Computation: Similar to existing methods, watermark scores for individual tokens are calculated through single token detection.
+Then, a small sliding window (i.e.
+$W=50$) is used to traverse the text to compute average scores within the window, serving as a smoothing operation.
+This results in a score list $s$ of length $N-W+1$, where $s_{i}$ represents the average watermark intensity from text token $x_{i}$ to $x_{i}+W$.
+(2) Anomaly Extraction: We design an anomaly extraction algorithm inspired by previous work in the field of style change detection (Zangerle et al., 2021) and intrinsic plagiarism detection (Manzoor et al., 2023).
+The mean score ($s_{\text{mean}}$) and top-k mean ($s_{\text{top-k-mean}}$) are calculated.
+Outliers are determined by:
+$\text{score}>s_{\text{mean}}+\frac{(s_{\text{top-k-mean}}-s_{\text{mean}})}{2}.$ (5)
+This is non-trivial because it ensures that the extracted suspicious watermarked regions likely cover the actual segments, with starting and ending deviations within a window size: when the sliding window falls entirely within the watermarked segment, the scores stabilize near $s_{\text{top-k-mean}}$; when the window falls completely outside, the scores stabilize below $s_{\text{mean}}$.
+Consequently, the extracted abnormal segment’s start and end points ($s^{\prime}$ and $e^{\prime}$) generally satisfy $s^{\prime}\in(s-W,s)$ and $e^{\prime}\in(e,e+W)$.
+Moreover, the use of $s_{\text{top-k-mean}}$ ensures good adaptability to various watermark strengths.
+The corresponding experiment results can be found at Section 6.5.
+(3) Fragment Connection: Adjacent outliers are connected with a predefined tolerance threshold, and segments shorter than a minimum length are filtered out, producing a list of segment indices.
 5.3 Local Traverse Detection
-After obtaining the localization results from the first step, the second phase (fine step), local traverse detection, conducts a more detailed verification.
-For each $(s^{\prime},e^{\prime})$ pair in the localization results, traverse inwards through the segments where the start point falls within $[s^{\prime},s^{\prime}+W)$ and the end point within $(e^{\prime}-W,e^{\prime}]$.
-Based on the previous analysis of the ranges of $s^{\prime}$ and $e^{\prime}$, this traversal has a high probability of reaching the gold index.
-Perform full-text detection sequentially on these segments, and select the statistic with the most significance to compare with the threshold.
-The pseudocode for the entire WaterSeeker algorithm can be found in Algorithm 3.
+After obtaining the coarse localization results, a fine-grained detection is performed by traversing segments within the predicted ranges.
+For each $(s^{\prime},e^{\prime})$ pair in the localization results, the algorithm examines segments with start points in $[s^{\prime},s^{\prime}+W)$ and end points in $(e^{\prime}-W,e^{\prime}]$.
+Based on the previous analysis, these ranges likely contain the true watermarked indices.
+Full-text detection is performed on these segments, and the most significant statistic is compared against a threshold.
+The complete WaterSeeker algorithm is presented in Algorithm 1.
+Algorithm 1 WaterSeeker Algorithm
+1:procedure Localization(tokens)
+2:     scores = SlidingWindow(tokens)
+3:     threshold = $s_{\text{mean}}+\frac{(s_{\text{top-k-mean}}-s_{\text{mean}})}{2}$
+4:     return ConnectOutliers(scores>threshold)
+5:end procedure
+6:procedure Detection(tokens, segs)
+7:     detected = []
+8:     for $(s^{\prime},e^{\prime})$ in segs do
+9:         best = $\infty$- ∞, idx = None
+10:         for $s\in[s^{\prime},s^{\prime}+W)$, $e\in(e^{\prime}-W,e^{\prime}]$ do
+11:              score=WatermarkScore(tokens[$s$:$e$])
+12:              if score > best then
+13:                  best = score
+14:                  idx = $(s,e)$
+15:              end if
+16:         end for
+17:         if best > threshold then
+18:              detected.append(idx)
+19:         end if
+20:     end for
+21:     return bool(detected), detected
+22:end procedure
+23:procedure WaterSeeker(tokens)
+24:     return Detection(tokens, Localization(tokens))
+25:end procedure
 5.4 Time Complexity Analysis
-Time Complexity of WaterSeeker: Suspicious Region Localization has a time complexity of $O(N)$.
-This step involves score list computation, anomaly detection, and fragment connection, each with a time complexity of $O(N)$, where $N$ is the text length.
-Local Traverse Detection has a time complexity of $O(W^{2})$.
-For each suspicious region, this step examines $W^{2}$ windows, where $W$ is the window size.
-The total time complexity of WaterSeeker is $O(N+W^{2})$.
+Time Complexity of WaterSeeker.
+WaterSeeker consists of two main components: (1) Suspicious Region Localization with $O(N)$ complexity, where $N$ is the text length, and (2) Local Traverse Detection with $O(W^{2})$ complexity, where $W$ is the window size.
+The total complexity is $O(N+W^{2})$.
 In practice, $W^{2}$ is typically kept lower than $N$, as a slightly larger window (i.e., $W=50$, detailed in Appendix E) suffices for a smooth and low-noise representation of the surrounding watermark intensity.
 Thus, the overall time complexity of WaterSeeker is $O(N)$.
-Lower Bound Complexity for the Problem: To detect watermarked segments in a long text, any algorithm must examine each token in the text at least once.
+Lower Bound Complexity for the Problem.
+To detect watermarked segments in a long text, any algorithm must examine each token in the text at least once.
 This requirement establishes a lower bound of $\Omega(N)$ for the time complexity of the problem, as at least one full pass through the text is necessary.
@@
 6 Experiment
-(a) KGW, $\delta=2.0$
-(b) KGW, $\delta=1.5$
-(c) KGW, $\delta=1.0$
-(d) Aar, temp=0.3
-(e) Aar, temp=0.2
-(f) Aar, temp=0.1
-Figure 3: This figure compares the detection performance of WaterSeeker and FLSW under varying watermark lengths and strengths, using Llama-2-7b as the generation model.
-Table 2: Average detection time per sample for various detection methods ($N\approx 10,000$, Unit: s).
-Full-text Detection WinMax FLSW-100 FLSW-200 FLSW-300 FLSW-400 WaterSeeker
-KGW 1.70 14.16 1.76 1.76 1.79 1.80 1.75
-Aar 0.54 2733.78 1.62 1.65 1.66 1.64 1.68
 6.1 Experiment Settings
 Watermarking Methods and Language Models: We selected two representative watermarking algorithms, KGW (Kirchenbauer et al., 2023) and Aar (Aaronson and Kirchner, 2022), each at three strength levels.
-KGW’s strength was set by the $\delta$ parameter (2.0=strong, 1.5=medium, 1.0=weak), while Aar’s strength used the temperature parameter (0.3=strong, 0.2=medium, 0.1=weak).
+KGW’s strength was set by the $\delta$ parameter (2.0=strong, 1.5=medium, 1.0=weak), while Aar’s strength used the temperature parameter (0.5=strong, 0.4=medium, 0.3=weak).
 We used Llama-2-7b (Touvron et al., 2023) and Mistral-7b (Jiang et al., 2024) as generation models.
@@
 For positive examples, one such segment was randomly inserted into each 10,000-token Wikipedia passage (Foundation, ).
+While most experiments used single-segment insertion, we also conducted experiments with multiple watermarked segments inserted into the same passage, with results reported in Table 3.
 Negative examples consist of unmodified 10,000-token Wikipedia corpus.
 Based on this procedure, four datasets were created, each containing 300 positive and 300 negative examples: KGW-Llama, KGW-Mistral, Aar-Llama, and Aar-Mistral.
-The specific distributions of watermark strengths and lengths are detailed in Appendix F.
-Baselines: As introduced in Section 4, Full-text Detection and WinMax (Kirchenbauer et al., 2024) are chosen, as well as Fix-Length Sliding Window method with $W$ of 100, 200, 300 and 400.
+Baselines: As introduced in Section 4, we selected Full-text Detection and WinMax (Kirchenbauer et al., 2024) with varying window size intervals, along with the Fixed-Length Sliding Window method using $W$ of 100, 200, 300, and 400.
 Hyper-parameters: The parameters related to WaterSeeker are as follows: $W=50,k=20$, with a tolerance for fragment connection set to 100.
-The parameter $\theta_{1}$ is set to 0.5, and $\theta_{2}$ is set to 1.5 for both algorithms.
 The threshold selection within the specified window is detailed in Appendix D.
 Notably, careful threshold selection is crucial for maintaining an acceptable false positive rate, as traversing long texts is prone to accumulating false positives.
-Evaluation-related parameters: $\theta$ mentioned in Section 3.2 is set to 0.5.
-Table 3: This table presents the contributions of the first step of WaterSeeker: Suspicious Segment Localization.
-It lists the average coverage of localization results relative to the ground truth segments under various watermark algorithms and watermark strengths, as well as the average offsets of the detected start and end indices.
-Metrics KGW Aar
-$\delta=2.0$ $\delta=1.5$ $\delta=1.0$ temp=0.3 temp=0.2 temp=0.1
-Average Coverage 0.992 0.984 0.953 0.996 0.991 0.977
-Average Offset (Start) 20.799 30.057 13.992 22.559 30.546 20.624
-Average Offset (End) 25.985 25.299 28.030 29.459 33.059 30.304
-Table 4: Comparison of detection performance with and without Local Traverse Detection across four datasets.
-Model Setting KGW Aar
-FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$ FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$
-Llama-2-7b w. local traverse 0.017 0.313 0.806 0.624 0.010 0.440 0.713 0.507
-w/o local traverse 0.003 0.390 0.756 0.520 0.000 0.450 0.710 0.462
-Mistral-7b w. local traverse 0.007 0.290 0.827 0.641 0.010 0.390 0.753 0.542
-w/o local traverse 0.003 0.380 0.763 0.534 0.003 0.413 0.738 0.490
-6.2 Detection Performance Analysis
-In Table 1, we present the detection performance of WaterSeeker alongside various baseline algorithms, evaluated on four datasets.
-Given that full-text detection lacks localization capabilities, its FNR is necessarily 1 under the evaluation rule requiring IoU $>\theta$ for successful detection.
-Therefore, we additionally report the FNR and corresponding F1 score for Full-text Detection without the IoU $>\theta$ constraint, highlighted in red.
-Additional performance results for WaterSeeker under various mixing ratios are presented in Appendix G.
-The results demonstrate that WaterSeeker achieves lower FNR and higher F1 score, significantly surpassing other watermark detection algorithms, including Full-text Detection, FLSW-100, FLSW-200, FLSW-300, and FLSW-400, while being comparable to WinMax.
-Based on the principle that WinMax is guaranteed to reach the gold index by evaluating all possible windows, it represents the upper bound of detection performance.
-However, subsequent experiments show that WinMax has extremely high time complexity (Section 6.4).
-Furthermore, the IoU results in the table reveal that WaterSeeker possesses a significantly higher localization capability compared to the baselines, again comparable to WinMax.
-6.3 Adaptability Analysis
-While the main experiment used four datasets with a mix of watermarked segments of varying strengths and lengths, this section compares the detection capabilities of WaterSeeker and the FLSW method across specific watermark strengths and lengths.
-Figure 3 shows that the performance of different detection algorithms follows a consistent trend: watermarked segments with higher strengths and longer lengths are more easily detected.
-FLSW performs well within a length range close to its window size but shows poor adaptability in other ranges due to its fixed-length nature, which leads to dilution effects when detecting longer or shorter watermarked segments.
-In contrast, WaterSeeker exhibits good adaptability across different length ranges by using anomaly extraction techniques and incorporates a top-k score mechanism to adapt to varying watermark strengths.
-6.4 Time Complexity Analysis
-Table 2 shows the average time spent per sample for WaterSeeker and other baseline algorithms during detection.
-WaterSeeker’s time is comparable to full-text detection and FLSW, and significantly lower than WinMax.
-Sections 4.2 and 5.4 analyze the theoretical time complexity of WinMax and WaterSeeker, respectively.
-This section offers a more comprehensive analysis based on experimental results, considering constant factors and investigating the significant disparity in WinMax’s execution time between KGW and Aar.
-Let $t_{1}$ be the time to compute a score in the score list, and $t_{2}$ be the time to compute the watermark detection statistic within a local window.
-WinMax’s time expenditure to detect a sample is $N\cdot t_{1}+N\cdot(W_{\max}-W_{\min})\cdot t_{2}$.
-For WaterSeeker, the localization step consists of score list computation, anomaly detection, and fragment connection, each with complexity $O(N)$.
-Denoting the time for the latter two sub-steps as $t_{3}\cdot N$, the total time for the first step is $N\cdot t_{1}+N\cdot t_{3}$.
-The local traverse detection step requires $W^{2}\cdot t_{2}$ time.
-Thus, WaterSeeker’s total time expenditure is $N\cdot t_{1}+N\cdot t_{3}+W^{2}\cdot t_{2}$.
-For the Aar (Aaronson and Kirchner, 2022) watermarking method, calculating the watermark detection statistic involves a complex GammaTransform function, resulting in a larger $t_{2}$.
-This leads to a significant time difference between WinMax and WaterSeeker due to the large disparity in coefficients preceding $t_{2}$.
-In this experiment, with $N=10,000,W_{\max}=400,W_{\min}=100,W^{2}=2,500$, the coefficient difference is a factor of 1,200.
-For the KGW (Kirchenbauer et al., 2023) watermarking method, calculating the z-score within a local window is relatively straightforward, reducing the time difference between WinMax and WaterSeeker.
-WaterSeeker is a general watermark detection method with consistently low time expenditure across various watermarking algorithms.
-In contrast, WinMax exhibits extremely high time costs when calculating the watermark detection statistic is complex, making it challenging for practical use.
+6.2 Results of Full-text Detection
+Table 1: Results of full-text detection methods.
+[…]
+KGW-Llama 0.000 0.983 0.033
+KGW-Mistral 0.000 0.993 0.013
+Aar-Llama 0.000 0.980 0.039
+Aar-Mistral 0.000 0.980 0.039
+Table 1 shows that full-text detection methods perform poorly across all four datasets, with an F1 score of less than 0.1.
+This indicates that full-text detection methods are totally ineffective for detecting watermarked segments in large documents.
+6.3 WaterSeeker Compared with WinMax
+Table 2: We evaluated the detection performance of WaterSeeker against various methods, including Full-text Detection, WinMax (Kirchenbauer et al., 2024), and FLSW.
+Results in this table utilized the Llama-2-7b model; results for Mistral-7b are presented in Appendix B.
+The metrics reported include false positive rate (FPR), false negative rate (FNR), F1 score, average Intersection over Union (IoU) between detected and ground truth segments, and processing time per sample.
+Best performances are highlighted in bold, while the second-best are underlined.
+Method KGW Aar
+[…]
+WinMax-1 0.017 0.193 0.885 0.713 1632.11 0.017 0.277 0.831 0.616 3615.42
+WinMax-50 0.017 0.220 0.868 0.672 34.31 0.007 0.307 0.816 0.577 72.13
+WinMax-100 0.013 0.237 0.859 0.632 17.16 0.003 0.330 0.800 0.554 35.34
+WinMax-200 0.010 0.273 0.834 0.547 9.12 0.003 0.363 0.776 0.486 18.38
+WaterSeeker 0.017 0.213 0.872 0.675 1.75 0.010 0.300 0.819 0.578 0.41
+FLSW-100 0.003 0.383 0.761 0.451 1.76 0.003 0.440 0.716 0.403 1.31
+FLSW-200 0.003 0.300 0.822 0.487 1.76 0.000 0.380 0.765 0.411 1.29
+FLSW-300 0.007 0.340 0.792 0.383 1.76 0.000 0.413 0.739 0.306 1.29
+FLSW-400 0.003 0.407 0.743 0.275 1.75 0.000 0.443 0.715 0.228 1.27
+Figure 3: A detailed comparison of WinMax and WaterSeeker regarding their detection performance and time cost, as well as their performance across varying text lengths.
+Main Experiment.
+From the data in Table 2, we can compare the detection and localization capabilities of WinMax with different window size intervals (1, 50, 100, 200) and WaterSeeker for watermarked segments, as well as the time costs for processing individual samples.
+WaterSeeker’s detection and localization performance is only slightly behind that of WinMax-1, while achieving a time savings of 1000 times.
+Given that WinMax evaluates all possible windows to ensure it reaches the gold index, it represents the upper bound of detection performance.
+However, as the window interval sizes for WinMax increase, the processing time decreases linearly with the interval size, yet it remains higher than that of WaterSeeker, while its detection and localization performance falls below that of WaterSeeker.
+Note: Table 2 demonstrates the results using Llama-2-7b, results for Mistral-7b can be found in Appendix B, which shows consistent trend.
+Balancing Performance and Time Cost.
+To clearly compare the balance between detection performance and time cost for WinMax and WaterSeeker, we collected additional data points for WinMax, as shown in Figure 3(a).
+Points further to the bottom right in the figure indicate a superior balance.
+It is evident that WaterSeeker is positioned clearly below and to the right of the curve formed by the WinMax data points, demonstrating a better balance between detection performance and time cost.
+Robustness against Varying Text Lengths.
+To further validate the robustness of watermark detection algorithms against varying document lengths (mixing ratios), we tested WaterSeeker and WinMax at $N=$ 500, 2000, 5000, and 10000, measuring the detection F1 score and time cost, as illustrated in Figure 3(b), (c).
+Both WaterSeeker and WinMax exhibit stable detection performance with changes in $N$.
+However, WinMax’s time cost increases at a higher rate than that of WaterSeeker, indicating its impracticality for large documents.
+Table 3: Comparison of detection and localization performance between WaterSeeker and FLSW with multiple segment insertion (one document containing three watermarked segments), including the time cost for processing each sample.
+Method KGW Aar
+[…]
+FLSW-100 0.000 0.177 0.903 0.433 1.76 0.013 0.147 0.914 0.434 1.31
+FLSW-200 0.000 0.110 0.941 0.475 1.75 0.007 0.093 0.948 0.461 1.29
+FLSW-300 0.003 0.130 0.929 0.399 1.76 0.007 0.120 0.933 0.369 1.29
+FLSW-400 0.003 0.153 0.916 0.314 1.76 0.003 0.150 0.917 0.288 1.27
+WaterSeeker 0.010 0.057 0.966 0.649 1.89 0.010 0.057 0.966 0.542 0.83
+WinMax’s Limitations for Multiple Segments.
+As shown in Equation 2, when multiple watermarked segments are inserted within the same document, WinMax cannot function properly.
+In contrast, WaterSeeker is able to adapt to this situation.
+The detection and localization performance for multiple watermarked segments can be found in the Table 3, comparing performance of WaterSeeker and FLSW.
+6.4 WaterSeeker Compared with FLSW
+Main Experiment.
+As shown in Table 2, although the time cost of FLSW is comparable to that of WaterSeeker, its detection performance is significantly inferior.
+This difference is due to FLSW’s fixed-length nature, which restricts its capability to utilize the gold index for detecting watermarked segments of varying lengths.
+Results for Mistral-7b is shown in Appendix B.
+Multiple Segments Detection.
+Table 3 presents the detection and localization results for documents containing three watermarked segments.
+The experimental setup mirrors that of the main experiment, utilizing Llama-2-7b as the generation model.
+The results indicate that as the number of inserted segments increases, it becomes easier to detect a watermarked segment (all methods show improved F1 scores).
+However, the Intersection over Union (IoU) did not exhibit significant changes.
+Notably, WaterSeeker continues to outperform the FLSW algorithm across all four configurations, consistent with the main experiment that included only one segment per sample.
+Further Analysis of Fix-length Nature.
+To better illustrate FLSW’s inability to adapt to watermarked segments of varying lengths, we selected two types of extreme examples from the main experiment dataset.
+As shown in Table 4, for segments with strong watermark intensity but short length, using larger window sizes such as FLSW-300 or FLSW-400 leads to the inclusion of many non-watermarked segments, which dilutes the watermark intensity and results in a significant drop in performance.
+Conversely, for segments with weak watermark intensity but longer length, using smaller window sizes like FLSW-100 or FLSW-200 results in an insufficient number of watermark tokens for detection, preventing the accumulation of intensity and adversely affecting the detection results.
+Table 4: Performance of WaterSeeker and FLSW in two types of examples: segments with strong watermark intensity but short length (length < 150, KGW $\delta=2.0$), and segments with weak watermark intensity but long length (length > 350, KGW $\delta=1.0$).
+TPR IoU
+Strong but Short
+FLSW-300 0.000 0.000
+FLSW-400 0.000 0.000
+WaterSeeker 0.667 0.572
+Weak but Long
+FLSW-100 0.375 0.150
+FLSW-200 0.625 0.447
+WaterSeeker 0.813 0.642
 6.5 Ablation Study
+Table 5: This table shows the contributions of the first stage of WaterSeeker: Suspicious Segment Localization.
+It lists the average coverage of localization results compared to ground truth segments for various watermark algorithms and strengths, along with the average offsets of the detected indices.
+Strength Avg. Cov. Avg. Off.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 6: Comparison of detection performance with and without Local Traverse Detection.
+Settings KGW Aar
+F1 IoU F1 IoU
+w. Traversal 0.872 0.675 0.819 0.578
+w/o Traversal 0.817 0.576 0.765 0.509
 We analyze the effectiveness of the two stages of WaterSeeker through an ablation study.
-The first stage, Suspicious Segment Localization, aims to achieve high coverage of the ground truth segments while keeping the start and end offsets within a specified window size.
-This ensures subsequent local traversals can reach the gold index.
-Table 3 shows that Step 1 achieves an average coverage exceeding 0.95 across various watermark algorithms and strengths, with average start and end offsets less than 50, remaining within the designated window size.
-Step 1 fulfills its purpose effectively.
-Local Traverse Detection performs a localized traversal based on the segments from Step 1, allowing for more refined verification within the window.
-Table 4 shows that across different LLMs and watermarking algorithms, Local Traverse consistently enhances detection F1 score and average IoU compared to directly applying full-text detection with the localization results, making it an indispensable component of WaterSeeker.
+Stage 1: Suspicious Region Localization.
+This stage aims to narrow down the detection target from a large document to a smaller region.
+The goal is to achieve high coverage of the ground truth segments while maintaining the start and end offsets within a specified window size.
+This ensures that subsequent local traversals can access the gold index.
+As shown in Table 5, Step 1 achieves an average coverage exceeding 0.9, with average start and end offsets remaining below $W$ across various watermark algorithms and strengths, demonstrating good adaptability.
+Stage 2: Local Traverse Detection.
+Local Traverse Detection performs a localized iteration based on the segments from Stage 1, allowing for more refined verification within the window.
+Table 6 shows that across different watermarking algorithms, Local Traverse consistently enhances detection F1 score and average IoU compared to directly applying detection with the localization results, making it an indispensable component of WaterSeeker.
+6.6 Robustness against Text Edit Attacks
+Figure 4: Robustness of WinMax, FLSW, and WaterSeeker against text edit attacks.
+The detection F1 score is reported for no attack, random word deletion attack (ratio = 0.3), and random word substitution attack (ratio = 0.3), along with the time cost for processing one sample.
+In this section, we further examine WaterSeeker’s robustness against text edit attacks, as watermarked segments generated by LLMs may be altered before integration into non-watermarked documents.
+Figure 4 illustrates the detection robustness of WaterSeeker, WinMax and FLSW against random word deletion (ratio=0.3) and substitution attacks (ratio=0.3, utilizing WordNet (Fellbaum, 1998) for synonym sets).
+The figure also contrasts these results with those obtained under no attack, along with the associated time costs.
+In this experiment, the KGW watermarking algorithm was employed, and Llama-2-7b is utilized as generation model.
+Since text editing can weaken the strength of the watermark, the watermark fragment intensity was set to a strong level under the "No attack" condition, specifically $\delta=2.0$.
+It can be observed from the figure that WaterSeeker achieves strong robustness against word deletion and word substitution attacks, with an F1 score exceeding 0.9.
+Compared with other baselines, only WinMax-1 perform slightly better than WaterSeeker, but at a significant cost in terms of time.
 7 Conclusion
@@
 Enhancing the sensitivity of WaterSeeker to detect shorter and weaker watermarks is an area for future improvement, which may involve refining the anomaly extraction algorithms or incorporating additional contextual analysis.
-Lastly, parameter selection, such as threshold settings, is crucial and can be challenging in different environments.
-Currently, parameters are manually tuned based on observations, so developing adaptive tuning methods, potentially using machine learning, would enhance WaterSeeker’s robustness and practicality across diverse scenarios.
+Lastly, parameter selection, including threshold settings to control false positives, is crucial and can be challenging in different environments.
+Stricter threshold controls can reduce the detection rate, necessitating adjustments based on the specific requirements of the actual settings.
 Appendix A Details of Representative Watermarking Algorithms
@@
 where $\Gamma$ is the Gamma Transformation function that converts the sum of correlation values to a p-value.
-Appendix B Pseudocode of WaterSeeker and Detection Baselines
-Pseudocode of WaterSeeker, WinMax and FLSW could be found in Algorithm 3, 1 and 2, respectively.
+Appendix B Supplementary Experimental Results Using Mistral-7b
+Table 7: We evaluated the detection performance of WaterSeeker against various methods, including Full-text Detection, WinMax (Kirchenbauer et al., 2024), and FLSW.
+Results in this table utilized Mistral-7b as the generation model.
+The metrics reported include false positive rate (FPR), false negative rate (FNR), F1 score, average Intersection over Union (IoU) between detected and ground truth segments, and processing time per sample.
+Best performances are highlighted in bold, while the second-best are underlined.
+Method KGW Aar
+[…]
+WinMax-1 0.010 0.243 0.857 0.641 1632.11 0.013 0.297 0.819 0.588 3615.42
+WinMax-50 0.007 0.270 0.841 0.613 34.31 0.007 0.320 0.806 0.560 72.13
+WinMax-100 0.003 0.283 0.833 0.588 17.16 0.007 0.333 0.797 0.531 35.34
+WinMax-200 0.003 0.340 0.794 0.511 9.12 0.003 0.347 0.789 0.501 18.38
+WaterSeeker 0.010 0.253 0.850 0.634 1.75 0.010 0.300 0.819 0.563 0.41
+FLSW-100 0.000 0.463 0.698 0.393 1.76 0.000 0.473 0.690 0.367 1.31
+FLSW-200 0.000 0.373 0.770 0.426 1.76 0.003 0.387 0.759 0.412 1.29
+FLSW-300 0.003 0.437 0.719 0.331 1.76 0.003 0.387 0.759 0.334 1.29
+FLSW-400 0.003 0.540 0.629 0.218 1.75 0.007 0.440 0.715 0.249 1.27
+Supplementary results for main experiment using Mistral-7b as generation model is shown in Table 7, showing consistent trend with Llama-2-7b.
+Appendix C Pseudocode of Detection Baselines
+Pseudocode of WinMax and FLSW could be found in Algorithm 2 and 3, respectively.
 Algorithm 2 WinMax Algorithm
-1:procedure WinMaxDetection(tokens, $W_{min}$, $W_{max}$, threshold)
+1:procedure WinMaxDetection(tokens, interval, threshold)
 2:     hasWatermark $\leftarrow$ False, indices $\leftarrow$ [ ]
 3:     maxStat $\leftarrow$ - $\infty$, bestIndex $\leftarrow$ None
-4:     for $W$in [$W_{min}$, $W_{max}$do
+4:     for $W$in [1, len(tokens)], step=interval do
 […]
-6:              stat $\leftarrow$ CalculateStatistics(tokens[$i:i+W$])
+6:              stat $\leftarrow$ WatermarkScore(tokens[$i:i+W$])
 7:              if stat $>$ maxStat then
@@
 […]
-5:         stat $\leftarrow$ CalculateStatistics(tokens[$i:i+W$])
+5:         stat $\leftarrow$ WatermarkScore(tokens[$i:i+W$])
 6:         if stat $>$ threshold then
@@
 13:end procedure
-Algorithm 3 WaterSeeker Algorithm
-1:procedure SuspiciousRegionLocalization(tokens, $W$, $k$, $\theta_{1}$, $\theta_{2}$)
-2:     Compute score list using sliding window
-3:     Calculate $s_{\text{mean}}$, $s_{\text{top-k-mean}}$, and $s_{\text{std}}$
-4:     Detect anomalies using threshold:
-5:     $s_{\text{mean}}+\max((s_{\text{top-k-mean}}-s_{\text{mean}})\cdot\theta_{1},s_{\text{std}}\cdot\theta_{2})$
-6:     Connect nearby outliers and filter short segments
-7:     return filteredSegments
-8:end procedure
-9:procedure LocalTraverseDetection(tokens, suspiciousRegions, $W$, threshold)
-10:     hasWatermark $\leftarrow$ False
-11:     indices $\leftarrow$ [ ]
-12:     for $(s^{\prime},e^{\prime})$ in suspiciousRegions do
-13:         maxScore = $\infty$- ∞, maxIndice = ()
-14:         for $s\in[s^{\prime},s^{\prime}+W)$ and $e\in(e^{\prime}-W,e^{\prime}]$ do
-15:              if DetectWatermark(tokens[$s:e$])$>$ maxScore then
-16:                  maxScore = DetectWatermark(tokens[$s:e$])
-17:                  maxIndice = $(s,e)$
-18:              end if
-19:         end for
-20:         if maxScore $>$ threshold then
-21:              hasWatermark $\leftarrow$ True
-22:              indices.append(maxIndice)
-23:         end if
-24:     end for
-25:     return hasWatermark, indices
-26:end procedure
-27:procedure WaterSeeker(tokens, $W$, $k$, $\theta_{1}$, $\theta_{2}$, threshold)
-28:     suspiciousRegions $\leftarrow$ SuspiciousRegionLocalization(tokens, $W$, $k$, $\theta_{1}$, $\theta_{2}$)
-29:     hasWatermark, indices $\leftarrow$ LocalTraverseDetection(tokens, suspiciousRegions, W, threshold)
-30:     return hasWatermark, indices
-31:end procedure
-Appendix C Detailed Proof for Aar
-C.1 Expected p-value of watermarked text
-In this sub-section, we aim to analyze how the expected p-value for watermarked text varies with changes in the detection window size $W$.
-Recall the p-value calculation formula:
-[…]
-where $S=\sum_{i=1}^{W}\log(\frac{1}{1-u_{i}})$, and $W$ is the window size.
-For watermarked tokens, $E[\log(\frac{1}{1-u_{i}})]=\mu_{1}$ , and for non-watermarked tokens, $E[\log(\frac{1}{1-u_{i}})]=\mu_{0}$, where $\mu_{1}>\mu_{0}$.
-The expectation of $S$ for different W values is as follows:
-  • When $W\leq L$: $E[S]=W\mu_{1}$
-  • When $W>L$: $E[S]=L\mu_{1}+(W-L)\mu_{0}$
-The expectation of p-value is:
-$E[\text{p-value}]=E[\Gamma(S,W,\text{loc}=0,\text{scale}=1)]$ (13)
-Since GammaTransform is non-linear, we cannot directly substitute $E[S]$. However, we can use Jensen’s inequality to obtain an approximation:
-$E[\Gamma(S,W,0,1)]\geq\Gamma(E[S],W,0,1)$ (14)
-Therefore, we can analyze $\Gamma(E[S],W,0,1)$ to obtain a lower bound for the expectation of p-value. Define function $f(W)=\Gamma(E[S],W,0,1)$:
-  • When $W\leq L$: $f(W)=\Gamma(W\mu_{1},W,0,1)$
-  • When $W>L$: $f(W)=\Gamma(L\mu_{1}+(W-L)\mu_{0},W,0,1)$
-Analyzing the behavior of $f(W)$:
-  • When $W\leq L$, both $E[S]$ and the shape parameter W increase as $W$ increases.
-  • When $W$ just exceeds $L$, the growth rate of $E[S]$ suddenly decreases (from $\mu_{1}$ to $\mu_{0}$), while the shape parameter W continues to increase linearly.
-Consider the behavior of $f(W)$ near $W=L$:
-  • When $W$ increases from $L-\epsilon$ to $L$, $E[S]$ increases by $\epsilon\mu_{1}$, and the shape parameter increases by $\epsilon$.
-  • When $W$ increases from $L$ to $L+\epsilon$, $E[S]$ increases by $\epsilon\mu_{0}$, and the shape parameter increases by $\epsilon$.
-Since $\mu_{1}>\mu_{0}$, at the point $W=L$, the rate of decrease of $f(W)$ suddenly slows down.
-This suggests that $f(W)$ is likely to reach its minimum value at $W=L$, or at a point very close to $L$.
-While this analysis does not strictly prove that the expectation of p-value is minimized exactly at $W=L$, it strongly suggests that the expectation of p-value is likely to reach its minimum value at or very near $W=L$ [^1] For a more rigorous proof, a deeper mathematical analysis of the GammaTransform function or numerical simulations would be necessary.
-C.2 Calculation of p-threshold to control false positive rate
-Similar to the analysis for KGW method, we also need to analyze the constraints on the p-value threshold $p^{*}$ when the false positive rate within the specified window is set to be lower than a target value $\alpha$.
-For non-watermarked text, $u_{i}\sim\text{Uniform}([0,1])$.
-Consequently, the test statistic $S$ follows a Gamma distribution: $S\sim\text{Gamma}(W,1)$, where $W$ is the shape parameter and 1 is the scale parameter.
-The p-value is calculated using the Gamma CDF:
-[…]
-where GammaCDF is the cumulative distribution function of the Gamma distribution with shape parameter $W$ and scale parameter 1.
-To achieve a false positive rate of $\alpha$, we need to set a threshold $p^{*}$ such that: $P(\text{p-value}<p^{*})=\alpha$.
-Given the definition of p-value, this is equivalent to: $P(1-\text{GammaCDF}(S,W,1)<p^{*})=\alpha$, which can be rewritten as: $P(S>\text{GammaInv}(1-p^{*},W,1))=\alpha$, where GammaInv is the inverse of the Gamma CDF.
-Since $S$ follows a $\text{Gamma}(W,1)$ distribution for non-watermarked text, we can express this as:
-$1-\text{GammaCDF}(\text{GammaInv}(1-p^{*},W,1),W,1)=\alpha.$ (16)
-Solving this equation for $p^{*}$, we get $p^{*}=\alpha$, which is also a constant value for different $W$.
 Appendix D Detail of Threshold Selection Within the Specified Window
@@
 D.2 Setting the threshold to achieve a target false positive rate $\alpha$
+KGW.
 For KGW, as analyzed in Section 5.1, when the window size is large, we can approximate using the Central Limit Theorem, resulting in $z^{*}=\Phi^{-1}(1-\alpha)$.
@@
 For segments shorter than 200, we use the binomial distribution and retrieve the corresponding threshold from the pre-computed dictionary.
-For Aar, as analyzed in Appendix C, $p^{*}=\alpha$, therefore set $p^{*}=10^{-6}$ for all lengths.
+Aar. For Aar, recall the p-value calculation formula:
+[…]
+where $S=\sum_{i=1}^{W}\log(\frac{1}{1-u_{i}})$, and $W$ is the window size.
+For non-watermarked text, $u_{i}\sim\text{Uniform}([0,1])$.
+Consequently, $S$ follows a Gamma distribution: $S\sim\text{Gamma}(W,1)$, where $W$ is the shape parameter and 1 is the scale parameter.
+Equation 17 is equivalent to:
+[…]
+where GammaCDF is the cumulative distribution function of the Gamma distribution with shape parameter $W$ and scale parameter 1.
+To achieve a false positive rate of $\alpha$, we need to set a threshold $p^{*}$ such that: $P(\text{p-value}<p^{*})=\alpha$.
+Given the definition of p-value, this is equivalent to: $P(1-\text{GammaCDF}(S,W,1)<p^{*})=\alpha$, which can be rewritten as: $P(S>\text{GammaInv}(1-p^{*},W,1))=\alpha$, where GammaInv is the inverse of the Gamma CDF.
+Since $S$ follows a $\text{Gamma}(W,1)$ distribution for non-watermarked text, we can express this as:
+$1-\text{GammaCDF}(\text{GammaInv}(1-p^{*},W,1),W,1)=\alpha.$ (19)
+Solving this equation for $p^{*}$, we get $p^{*}=\alpha$, which is also a constant value for different $W$.
 Appendix E Impact of Window Size on Watermark Intensity Calculation
@@
 The analysis encompasses the ground truth segment and 500 tokens on either side.
-Figures 4 and 5 illustrate the results for the KGW and Aar algorithms, respectively.
+Figures 5 and 6 illustrate the results for the KGW and Aar algorithms, respectively.
 The intensity curves reveal that small window sizes, particularly $W\leq 10$, introduce significant fluctuations.
@@
 Consequently, we adopt $W=50$ for our main experiments.
-Appendix F Sample Distribution in the Main Experiment
-It can be observed from Figure 6 that all four datasets include samples of watermarked segments with varying intensities and lengths, demonstrating the comprehensiveness and fairness of the dataset construction.
-(a) KGW, Llama
-(b) Aar, Llama
-(c) KGW, Mistral
-(d) Aar, Mistral
-Figure 6: Distribution of sample numbers for the four datasets involved in the main experiment.
-Appendix G Further Experimental Results under Different Mixing Ratios
-In main experiment, 100-400 watermarked tokens are inserted into non-watermarked text of 10,000 tokens, resulting in watermarked text mixing ratios of 1% to 4%.
-This section expands the evaluation by testing the performance of various methods across a broader range of mixing ratios, achieved by adjusting the length of the non-watermarked text.
-We deliberately avoid adjusting the mixing ratio by inserting multiple watermark segments into the non-watermarked text, as the WinMax method can only return the highest-scoring segment and is not designed for multi-segment insertion scenarios.
-Additionally, we do not use fixed-length watermark segments because in real-world detection scenarios, the length of the watermark segment is typically unknown.
-Moreover, if the watermark length were fixed and known, a simple Fixed-length Sliding Window detection method (with the window size equal to the watermark segment length) would suffice to solve the problem.
-We provided further experimental results under three different settings:
-  1. 1.
-    100-400 watermarked tokens are inserted into 5,000 non-watermarked tokens, with a mixing ratio of 2% (100 / 5100) -7.4% (400 / 5400). The results are demonstrated in Table 6.
-  2. 2.
-    100-400 watermarked tokens are inserted into 2,000 non-watermarked tokens, with a mixing ratio of 4.8% (100 / 2100) -16.7% (400 / 2400). The results are demonstrated in Table 7.
-  3. 3.
-    100-400 watermarked tokens are inserted into 500 non-watermarked tokens, with a mixing ratio of 16.7% (100 / 600) -44.4% (400 / 900). The results are demonstrated in Table 8.
-Table 6: Comparison of WaterSeeker with other baselines under the mixing ratio of 2%-7.4%.
-Model Method KGW Aar
-FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$ FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$
-Llama-2-7b Full-text Detection 0.000 1.000 [0.990] 0.000 [0.020] 0.001 0.000 1.000 [0.957] 0.000 [0.083] 0.002
-WinMax 0.003 0.317 0.810 0.621 0.010 0.380 0.761 0.565
-FLSW-100 0.000 0.520 0.649 0.413 0.003 0.523 0.644 0.375
-FLSW-200 0.000 0.457 0.704 0.404 0.000 0.560 0.611 0.350
-FLSW-300 0.003 0.663 0.502 0.297 0.000 0.767 0.378 0.283
-FLSW-400 0.003 0.867 0.235 0.230 0.000 0.920 0.148 0.205
-WaterSeeker(Ours) 0.003 0.357 0.781 0.570 0.010 0.407 0.740 0.536
-Mistral-7b Full-text Detection 0.000 1.000 [0.990] 0.000 [0.020] 0.000 0.000 1.000 [0.937] 0.000 [0.119] 0.004
-WinMax 0.003 0.227 0.871 0.700 0.010 0.370 0.768 0.570
-FLSW-100 0.000 0.503 0.664 0.443 0.000 0.543 0.627 0.364
-FLSW-200 0.000 0.350 0.788 0.497 0.003 0.540 0.629 0.362
-FLSW-300 0.000 0.527 0.643 0.392 0.003 0.763 0.382 0.269
-FLSW-400 0.000 0.783 0.357 0.281 0.003 0.910 0.165 0.203
-WaterSeeker(Ours) 0.000 0.243 0.861 0.676 0.010 0.383 0.758 0.548
-Table 7: Comparison of WaterSeeker with other baselines under the mixing ratio of 4.8%-16.7%.
-Model Method KGW Aar
-FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$ FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$
-Llama-2-7b Full-text Detection 0.000 1.000 [0.967] 0.000 [0.065] 0.005 0.000 1.000 [0.873] 0.000 [0.225] 0.018
-WinMax 0.003 0.257 0.851 0.669 0.003 0.413 0.738 0.533
-FLSW-100 0.000 0.473 0.690 0.440 0.003 0.570 0.600 0.350
-FLSW-200 0.000 0.420 0.734 0.446 0.000 0.607 0.565 0.327
-FLSW-300 0.003 0.630 0.539 0.339 0.000 0.777 0.365 0.257
-FLSW-400 0.003 0.850 0.260 0.246 0.000 0.883 0.209 0.200
-WaterSeeker(Ours) 0.003 0.287 0.831 0.642 0.003 0.443 0.714 0.501
-Mistral-7b Full-text Detection 0.000 1.000 [0.980] 0.000 [0.039] 0.003 0.000 1.000 [0.870] 0.000 [0.230] 0.017
-WinMax 0.000 0.220 0.876 0.710 0.003 0.353 0.783 0.590
-FLSW-100 0.000 0.470 0.693 0.447 0.000 0.487 0.678 0.415
-FLSW-200 0.000 0.360 0.780 0.497 0.003 0.500 0.665 0.371
-FLSW-300 0.000 0.533 0.636 0.386 0.003 0.760 0.386 0.285
-FLSW-400 0.000 0.747 0.404 0.288 0.003 0.887 0.203 0.212
-WaterSeeker(Ours) 0.000 0.227 0.872 0.701 0.003 0.363 0.776 0.574
-Table 8: Comparison of WaterSeeker with other baselines under the mixing ratio of 16.7%-44.4%.
-Model Method KGW Aar
-FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$ FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$
-Llama-2-7b Full-text Detection 0.000 1.000 [0.787] 0.000 [0.352] 0.081 0.000 1.000 [0.693] 0.000 [0.469] 0.109
-WinMax 0.000 0.247 0.859 0.685 0.000 0.373 0.770 0.568
-FLSW-100 0.000 0.470 0.693 0.445 0.000 0.557 0.614 0.380
-FLSW-200 0.000 0.383 0.763 0.461 0.000 0.553 0.618 0.340
-FLSW-300 0.000 0.553 0.618 0.360 0.000 0.690 0.473 0.298
-FLSW-400 0.000 0.740 0.413 0.265 0.000 0.837 0.281 0.231
-WaterSeeker(Ours) 0.000 0.277 0.839 0.647 0.000 0.410 0.742 0.522
-Mistral-7b Full-text Detection 0.000 1.000 [0.847] 0.000 [0.266] 0.061 0.000 1.000 [0.697] 0.000 [0.465] 0.112
-WinMax 0.000 0.270 0.844 0.652 0.000 0.350 0.788 0.588
-FLSW-100 0.000 0.510 0.658 0.432 0.000 0.530 0.639 0.393
-FLSW-200 0.000 0.407 0.745 0.442 0.000 0.510 0.658 0.369
-FLSW-300 0.000 0.570 0.601 0.352 0.000 0.690 0.473 0.294
-FLSW-400 0.000 0.760 0.387 0.243 0.000 0.827 0.295 0.246
-WaterSeeker(Ours) 0.000 0.303 0.821 0.612 0.000 0.407 0.745 0.543
-Experimental results demonstrate that WaterSeeker exhibits high robustness across various mixing ratios, showing consistent performance trends across different settings.
-Specifically, WaterSeeker significantly outperforms full-text detection and FLSW methods in both detection accuracy and localization effectiveness, while achieving comparable performance to the upper bound method, WinMax.
-Appendix H Further Application Scenario
-Figure 7: An illustration of WaterSeeker applied in AI detection system.
-WaterSeeker not only provides accurate detection results, but its localization capabilities also contribute to building a more transparent and interpretable AI detection system.
-As illustrated in Figure 7, the AI detection system powered by WaterSeeker receives documents for analysis and outputs three components: (1) Detection Result, which indicates whether AI assistance was utilized; (2) Suspicious AI-generated Segments, which highlights segments identified as potentially AI-generated; and (3) AI Ratio, which displays the proportion of content attributed to AI.
-For example, in the context of academic integrity, highlighting suspicious segments provides actionable evidence for assessments and appeals.
-Additionally, reporting the AI ratio allows for more flexible establishment of academic misconduct standards.